### v1

# A Comunication Framework for Compositional Generation

Abstract
Compositionality and compositional generalization—the ability to understand novel combinations of known concepts—are central characteristics of human language and are hypothesized to be essential for human cognition. In machine learning, the emergence of this property has been studied in a communication game setting, where independent agents (a sender and a receiver) converge to a shared encoding policy from a set of states to a space of discrete messages, where the receiver can correctly reconstruct the states observed by the sender using only the sender’s messages. The use of communication games in generation tasks is still largely unexplored, with recent methods for compositional generation focusing mainly on the use of supervised guidance (either through class labels or text). In this work, we take the first steps to fill this gap, and we present a self-supervised generative communication game-based framework for creating compositional encodings in learned representations from pre-trained encoder-decoder models. In an Iterated Learning (IL) protocol involving a sender and a receiver, we apply alternating pressures for compression and diversity of encoded discrete messages, so that the protocol converges to an efficient but unambiguous encoding. Approximate message entropy regularization is used to favor compositional encodings. Our framework is based on rigorous justifications and proofs of defining and balancing the concepts of Eficiency, Unambiguity and Non-Holisticity in encoding. We test our method on the compositional image dataset Shapes3D, demonstrating robust performance in both reconstruction and compositionality metrics, surpassing other tested discrete message frameworks.
Figure 1: General concept of the proposed framework. In a communication game a sender $S$ and receiver $R$ jointly aim to encode the observed latent states as symbolic messages. Useful length $\ell$ gives the minimum length of an initial segment of a $S$-encoded message, after which its $R$-reconstruction no longer improves. Above we show an example of images generated from increasing initial segments of a message, and in this case $\ell(x)=6$. Below we illustrate a typical alternation between interaction and imitation phases in IL: arrow tips indicate latent space $R$-encodings of increasing initial submessages of lenghts $1,\dots,\ell(x)$ of $S$-encoded messages.
Machine Learning, ICML
1 Introduction
Humans have the innate ability to segment and reimagine cognitive and perceptual representations, by modifying their components and characteristics, much like how we compose words into complex sentences to describe and understand the world around us (Arun, 2022).
One of the main challenges in modern Machine Learning is to faithfully capture this hability, also known as compositional generalization. In particular, the ability to imagine new combinations of known components translates directly into Compositionality in Generative models.
Compositional generation is directly relevant to the prominent challenges of learning long-tailed data distributions, in which models tend to mode collapse due to underrepresented classes (Saxena & Cao, 2021; Qin et al., 2023). It is often the case for long-tail distributions that several variable features of the data can be compositionally modified, producing a large combinatorial ensemble of different modes (Wiedemer et al., 2023), and thus compositional capabilities of a learning model will directly transfer to strong advances in Out Of Distribution (OOD) generalization (Liu et al., 2021; Abbasi et al., 2024).
The quest for efficient compositional priors for machine learning models has long been one of the main challenges of artificial intelligence. We should distinguish between several possible meanings for compositional generation:
  1. 1.
    Efficient generative transduction includes classical text-to-image generation tasks in which compositional dependencies are required for accurate OOD generation (Liu et al., 2022b; Leivada et al., 2023).
  2. 2.
    Controllable image generation requires to use, or detect and use, disentangled compositional features which will be controlled for correct generation (Nie et al., 2021; Casanova et al., 2023; Okawa et al., 2024)
  3. 3.
    Faithful generation of multimodal statistical distributions with latent compositional structure assumes that underlying multimodality of black-box target distributions, there is a compositional structure of the data. Here, no attempt at direct capture (or disentanglement) of this structure is made, and the focus is on imposing architecture priors to enhance the faithfulness of representation (Chowdhury et al., 2021; Latheef et al., 2024; Denamganai & Walker, 2020; Dessì et al., 2021).
We note that while the above three directions all require tackling the challenge of compositionality in different forms, cases 1 and 2 face the further difficulty of complex choice of metrics for testing their success. Here, we focus on direction 3 as it addresses a more fundamental question, in which generation accuracy objectively quantifies the success of capturing the full compositional complexity if we use a compositionally generated dataset. This relatively simpler setting allows us to focus precisely on the challenges of effective learning of compositional priors, leaving the extensions to more complex goals for future works.
Recent works have found a connection between simpler representations and compositionality (Elmoznino et al., 2024; Ren et al., 2023; Ren & Sutherland, 2024). Specifically, (Ren & Sutherland, 2024) found that compositional mappings are the simplest (smallest upper bound on Kolmogorov complexity) bijections between the underlying generating factors of a signal and the representation space.
Furthermore, as an independent benefit, from foundational works such as (Rissanen, 1978), penalizing the Kolmogorov Complexity (KC) of descriptions has been viewed as a computational version of Occam’s razor for learning models, inducing a kind of regularization leading to efficiency improvement for general learning frameworks (Hochreiter & Schmidhuber, 1997a; Hinton & Van Camp, 1993). In practice, indirect methods have to be adopted, because KC is impractical to compute, and furthermore in our case the coding underlying the compositional structure of our data is not assumed to be accessible. A promising approach has been the IL framework (Kirby, 2001), which implements active agents to learn a mapping from a meaning space into a signal space, modelled e.g. as a Definite Clause Grammar.
The IL mechanism can be modeled as a communication channel bottleneck (Havrylov & Titov, 2017; Kharitonov et al., 2020; Lerch & Sims, 2019; Resnick et al., 2019), with its success in KC regularization being justified via rate distortion theory (Tishby et al., 2000), but the main motivation for its validity comes from the analogy to the theory of how human and general languages are formed. Here the basic framework stems from sociolinguistic theory (Kirby & Hurford, 2002; Kirby et al., 2015; Blume et al., 1993; Bhui et al., 2021; Norris & Kalm, 2021; Rita et al., 2022; Lai, 2024), as well as from game theory formulations via the so-called referential signaling games introduced by Lewis (Lewis, 1969; Huttegger et al., 2010a; Huttegger & Zollman, 2011; Huttegger et al., 2010b; Barrett & Gabriel, 2021; Huttegger et al., 2014).
While, as mentioned above, several works show that IL allows for simplicity- or KC-based regularization (Mordatch & Abbeel, 2018; Li & Bowling, 2019b; Korbak et al., 2020; Ren et al., 2023, 2020; Wärneryd, 1993; Zhou et al., 2022; Li & Bowling, 2019a; Ren et al., 2020), these approaches are often built with language-based, transduction (Raj et al., 2017) or symbolic (Havrylov & Titov, 2017; Mathy & Feldman, 2012; Lazaridou et al., 2018) data in mind, based on their origin in the theory of language optimization (Foerster et al., 2016), attempts for vision have been focussed on notions of compositionality as in points 1 and 2 from our description (Choi et al., 2018; Nassar et al., 2018; Liu et al., 2022a; Devaraj et al., 2020; Latheef et al., 2024; Vani et al., 2021), and have shown limited success on the datasets which we work on (Schott et al., 2022). This latter work is one of the main motivations of this paper.
1.1 Related work and main contributions
As mentioned before, the use of IL for fitting compositional multimodal statistical distributions is underexplored. Related works such as (Chowdhury et al., 2021; Denamganai & Walker, 2020) feature a sender and receiver but do not develop a symbolic bottleneck that implies KC regularization as ours. Further works like (Dessì et al., 2021) focus on other aspects within a somehow related IL framework for image generation but do not tackle the bottleneck structure in depth. As far as we know, the only existing attempt to use emergent language models in image generation tasks is (Latheef et al., 2024), with which our work shares the choice of a comunication bottleneck, used for long-tailed image generation of brain activity patterns. This model, however, does not make use of the IL framework.
Our main contributions are the following:
  • We propose a joint implementation of the Lewis Reconstruction Game and IL framework, which to our best knowledge is the first instance of such framework that takes into account the peculiarities of self-supervised generation tasks.
  • We introduce a novel regularization technique to favor exploration and diversity in encoding, with mathematically justified tunable biases towards compositionality and encoding efficiency.
  • We accompany the above by a rigorous mathematical framework (developed in the appendices), that defines the setting and objectives of compositional generation, and justifies our architecture choices.
2 Background
2.1 Lewis reconstruction game
The Lewis reconstruction game (Rita et al., 2022) is a special case of the Lewis Signaling Game (LSG) (Lewis, 1969) framework, in which two agents cooperate to reconstruct an observed object. Specifically, a sender agent parameterized by $\varphi$ observes an object $x$ from the set of objects $\mathcal{X}$ and produces a message $m\in\mathcal{M}$, where $\mathcal{M}$ is the set of all posible strings formed by concatenating symbols drawn from a finite vocabulary $V$. A receiver agent parameterized by $\omega$ observes $m$ and produces a reconsruction $x^{\prime}$. Both models are optimized to minimize a reconstruction loss $\mathcal{L}_{R}(x,x^{\prime})$.
Importantly, in the original LSG, the receiver is tasked with predicting $x$ from a finite set of states, whereas in this formulation, the problem becomes a regression task, and $\mathcal{X}$ can be a continuous space.
2.2 Iterated Learning
The IL model of cultural evolution (Kirby, 2001) proposes to emulate the emergence of language through the interactions of adult and new learning agents. In the original formulation, a space of pairs of signal + meaning is randomly generated, and in each iteration, a learning agent is partially exposed to the language learned by the adult agent in the previous iteration. The learning agent becomes the adult agent for the next iteration, and transmits a subset of its learned ”language”. By iterating this process, the initial random language gains structure and observable properties, such as compositionality.
This framework has been widely extended to include an additional task (Li & Bowling, 2019a; Ren et al., 2023; Zhou et al., 2022), instead of simply reshaping the language. Under this formulation, the training regime can be separated into two parts: Interaction (Int), where an agent is optimized to solve the additional task, and Imitation (Im), where a newly initialized agent ”learns” from the previous agent through a shared learned language.
2.3 Compositional data generation assumptions
In a formalism similar to (Ren et al., 2023), we first model the images by a random variable $X=\mathsf{GenX}(\mathbf{G},\mathbf{O})$, in which data $X$ are created via deterministic function $\mathsf{GenX}$ from a set of generating factors $\mathbf{G}$, and a source of randomness $\mathbf{O}$. We have access to samples from $X$ but not to $\mathsf{GenX},\mathbf{G},\mathbf{O}$, and the goal is to reproduce the distribution of $X$, i.e., generate a random variable $\widehat{X}$ that is close to $X$ in a natural metric $\mathcal{L}(X,\widehat{X})$.
We restrict to the case that $\mathbf{G}$ has a compositional structure, e.g. $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which $G_{i}$ represent independent characteristics having finitely many possible values (as in the dataset of Sec. 5.1 for example). In this case, an efficient way to reproduce $X$ is to impose that $\widehat{X}$ is similarly generated as $\widehat{X}=\mathsf{GenX^{\prime}}(\mathbf{G^{\prime}},\epsilon)$ using a source of randomness $\epsilon$, where generating factors $\mathbf{G^{\prime}}$ should be a learned efficient encoding of $\mathbf{G}$. In fact we may argue, similar to (Ren et al., 2023), that if $\mathbf{G^{\prime}}$ is a compressed representation, then it tends to have low Kolmogorov complexity and to be isomorphic to the actual $\mathbf{G}$, leading to compositional generation of $\widehat{X}$. The arrow $X\to\mathbf{G^{\prime}}$ in the diagram below indicates the fact that $X$ is the datum from which elements from $\mathbf{G^{\prime}}$ have to be deduced for generating $\widehat{X}$, since $X$ is all we have access to.
$\mathbf{G}$$\mathbf{O}$$X$$\mathbf{G^{\prime}}$$\epsilon$$\widehat{X}$$\approx$ iso$\mathcal{L}(X,\widehat{X})$unknown
Here we focus on the case that $\mathbf{G}$ has a very large number of classes, in which case the main novel difficulty is to build a framework for finding the optimized compositional encoding $\mathbf{G^{\prime}}$ from observation of $X$ only. In order to avoid technical complications and confounding factors for our conclusions, we restrict to the simplified case that $\mathsf{GenX}$ generates a single element for each class in $\mathbf{G}$, in which case we can limit ourselves to the case of trivial noise sources $\mathbf{O},\epsilon$, and leave the challenges towards extension of the framework to cases of general $\mathbf{O},\epsilon$ to future study. Thus we assume that
$X=\mathsf{GenX}(\mathbf{G}),\quad\widehat{X}=\mathsf{GenX^{\prime}}(\mathbf{G^{\prime}}),$ (1)
[…]
In (1) the randomness of generated $X$ disappears, and effectively we work with a deterministic dataset $\mathcal{D}=\mathsf{Im}(\mathsf{GenX})\subseteq\mathcal{X}$ (where $\mathcal{X}$ is an ambient space, e.g. a ball in $\mathbb{R}^{D}$ for some large $D$) with one element for each class in $\mathbf{G}$, which we partition into $\mathcal{D}=\mathcal{D}_{train}\cup\mathcal{D}_{test}$ for training and testing. The compositional nature of $\mathbf{G}$ facilitates reconstruction up to isomorphism from just $\mathcal{D}_{train}$, a fact that is not guaranteed under general encodings (see §A for a formal treatment and Section 3.4 for discussion).
3 Proposed architecture
Figure 2: Overview of our proposed architecture: (Int) Sender $S_{\phi^{t}}$ and receiver $R_{\omega^{t}}$ are jointly trained to minimize the reconstruction error between the true state $x$ and predictions $\{x^{\prime}\}$ by encoding $x$ as a message $m^{t}_{x}$. (Im) The next sender $S_{\phi^{t+1}}$ is initialized to find a protocol that matches the final predicted state of $R_{\omega^{t}}(m^{t}_{x})$ and maximizes the pairwise distance between messages.
3.1 The signaling game approach
Our search space for $\mathbf{G^{\prime}}$ will consist of messages of length $\leq C$ consisting of words from a vocabulary $V$, i.e., we assume that $\mathbf{G^{\prime}}\in\mathcal{M}:=V^{C}$. The objective of finding the optimal encodings $\mathbf{G^{\prime}}$ is modeled as a signaling game: the map $S:\mathcal{X}\to\mathcal{M}$ plays the role of a sender that for each signal $x\in\mathcal{X}$ will produce some message $m\in V^{C}$, and $\mathsf{GenX^{\prime}}$ will be the restriction to $S(\mathcal{D})\subseteq\mathcal{M}$ of a suitable map $R:\mathcal{M}\to\mathcal{X}$, which represents a receiver that decodes the sender’s messages. The following diagram summarizes our notations:
$\mathcal{G}=\{\mathbf{G}\}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{D}=\mathsf{GenX}(\mathcal{G})\subseteq\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}=V^{C}\stackrel{{\scriptstyle R}}{{\longrightarrow}}\mathcal{X}$
Sender and receiver will be parametrized by recurrent neural networks, and have the common goal to find optimal signaling strategy, as measured by $\mathcal{L}(\cdot,\cdot)$.
In the context of the signaling game, the sender $S$ and receiver $R$ must establish a protocol that satisfies the following properties:
  1. (UE)
    Unambiguous Encoding: $S$ must produce an encoding of $x\in\mathcal{D}$ as messages $m\in\mathcal{M}$ that is unambiguous, i.e., injective. Specifically, states generated from distinct combinations of factors in $\mathbf{G}$ must be mapped to distinct messages by $S$.
  2. (EE)
    Efficient Encoding: The encoding of $m=S(x)$ should be such that $R$ is able to reconstruct any state $x\in\mathcal{D}$ by interpreting as few symbols from the message $m$ as possible. This property promotes efficiency in communication.
3.2 Unambiguity
As proposed by (Ren et al., 2023), compositionality emerges when $\mathbf{G}$ and $\mathbf{G^{\prime}}$ are isomorphic. In particular the map $\mathbf{G}\mapsto x\mapsto\mathbf{G^{\prime}}$ has to be injective. We assume the mapping $\mathbf{G}\mapsto x$ to be injective, thus the above composition being injective is equivalent to $x\mapsto\mathbf{G^{\prime}}$ being injective, which is condition UE.
However, achieving unambiguity is non-trivial. LSGs are known to exhibit two types of equilibria:
  • Separating Equilibria, where $S$ and $R$ coordinate to encode and decode messages that perfectly differentiate every state (UE).
  • Pooling Equilibria, where $S$ or $R$ associates multiple states with the same message.
Without appropriate incentive mechanisms, LSGs tend to converge to pooling equilibria due to their prevalence over separating equilibria (Huttegger et al., 2010b), a problem which (as shown in §D in an illustrative example) is exacerbated unless $|\mathcal{D}_{train}|$ remains relatively large.
This challenge is exacerbated in continuous state spaces. For instance, if two state vectors $x_{1}=\mathsf{GenX}(\mathbf{G}_{1}),$ and $x_{2}=\mathsf{GenX}(\mathbf{G}_{2})$, generated from $\mathbf{G}_{1},\mathbf{G}_{2}\in\mathcal{G}$, are sufficiently close (i.e., $d(x_{1},x_{2})<\delta$ for small $\delta$), typical loss functions give little incentive for $S$ to separate these states in $\mathcal{M}$: the best-case loss improvement that can occur when passing from coding $x_{1},x_{2}$ as the same message, to coding them as different messages, is bounded by $\sup_{x\in\mathcal{X}}|\mathcal{L}(x_{1},x)-\mathcal{L}(x_{2},x)|$, a value which tends to zero as $\delta\to 0$ in most situations of interest (e.g. whenever $\mathcal{L}$ is continuous and $\mathcal{D}\subset\mathcal{X}$ is bounded).
3.3 Efficiency
While EE is not strictly necessary for the theoretical soundness of the framework, it is a practical way to encourage generalization. This aligns with the information bottleneck principle and feature compression (Tishby et al., 2000; Kawaguchi et al., 2023). Additionally, the Kolmogorov complexity of the mapping $\mathbf{G^{\prime}}\mapsto\hat{X}$ depends on both the complexity of the model and the complexity of the inputs. Simplifying the message space reduces the complexity of the decoding scheme, facilitating learning and generalization.
3.4 Avoiding Holistic Encodings
A holistic encoding is an injective mapping $\mathcal{G}\to\mathcal{M}$ that encodes $\mathbf{G}$ into $\mathbf{G^{\prime}}$ but disregards the underlying characteristics of the elements (Kirby & Hurford, 2002) (see §A for a precise definition). For example, a holistic encoding might injectively enumerate each element $\mathbf{G}\in\mathcal{G}$ assigning to it a natural number $n_{\mathbf{G}}\in\mathbb{N}$. Since these mappings fail to reflect the structure of $\mathbf{G}$, they will fail to generalize well to unseen combinations of generating factors. Thus it is important to set up incentives for our communication game so that besides UE and EE it also achieves the following condition, which is favorable for compositional generalization:
  1. (NHE)
    Non Holistic Encodings: $S$ must avoid creating encodings that are holistic.
We assume that most combinations of the underlying factors are represented in some element of the training set $\mathcal{D}_{train}\subset\mathcal{D}$. This is reasonable to assume, even while assuming that $|\mathcal{D}_{train}|/|\mathcal{D}|$ is small, i.e., only a small portion of possible factor combinations are present (see §A.1 and in particular Thm. A.3 for a proof). In this case, compositional encodings could be deduced from $\mathcal{D}_{train}$ only. However, a naive optimization of efficiency objective EE on $\mathcal{D}_{train}$ necessarily leads to holistic encodings, a phenomenon that we illustrate in Example 3.1. Furthermore, in the large data $|\mathcal{D}|\gg 1$ regime holistic encodings have overwhelmingly higher Kolmogorov complexity than compositional ones, leading to poor generalization, we cannot guarantee UE and EE to generalize to $\mathcal{D}$. This is why enforcing NHE becomes necessary (for proofs of these claims and a formal discussion, see §B).
Example 3.1.
Consider a scenario with $k$ binary generating factors and a training set containing $|\mathcal{D}_{train}|=\frac{2^{k}}{8}=|\mathcal{D}|/8$ random samples from $\mathcal{D}$. The minimum number of bits required for a compositional mapping $\mathcal{G}\to\mathcal{M}$ is $\log 2^{k}=k$, whereas the minimum for a holistic mapping is $\log\frac{2^{k}}{8}=k-3$. Here, achieving EE solely over the training set is guaranteed to produce a holistic encoding that fails to generalize.
We then hypothesize that in order to favor NHE an effective heuristic is to maximize diversity (allowing larger ”useful lengths”) of encoded messages $\mathbf{G^{\prime}}$ (see §B).
4 Proposed architecture
Given a data distribution $\mathcal{X}$, a finite set of symbols $V$ and a fixed channel capacity $C$, we will define $\mathcal{M}=[V]^{C}$ as the set of all posible combinations of symbols in $V$ of length $C$, and sender and receiver $S$ and $R$ as learneable neural networks with parameters $\phi$ and $\omega$, respectively.
Following the IL framework (Kirby, 2001), we separate learning between an interaction phase and an imitation phase.
4.1 Interaction (Int)
During the Int phase, we jointly train $S$ and $R$ to optimize the reconstruction of $X$. To add compression pressure and enforce hierarchical representations and EE (Latheef et al., 2024), we include the following achitecture choices (i) we implement $S,R$ as recurrent neural networks, to further enforce the dependence of each symbol on the preceding ones; (ii) rather than optimizing the protocol to minimize the difference between the true state and the prediction given by the receiver after viewing the full message (Rita et al., 2022), the receiver makes a prediction with each new symbol from the message, and we minimize the average difference between the predicted states and the true states. For more explicit compression presure, we add a hyperparameter $\lambda\in[0,1]$. The Int phase loss becomes:
$\mathcal{L}^{int}_{\omega,\phi}:=\mathbb{E}_{x\sim\mathcal{X}}\left[\frac{1}{C}\sum_{i}^{C}\lambda^{i}d_{\mathcal{X}}(x,R_{\omega^{t}}(S_{\phi^{t}}(x)_{[i]}))\right]$ (2)
Where $[i]$ refer to the first $i$ symbols of a message and $d_{\mathcal{X}}$ will be a suitable distance or discrepancy measure over $\mathcal{X}$.
Remark 4.1 ($\mathcal{L}^{int}_{\omega,\phi}$ is a natural loss for compositional generative tasks).
If $x,x^{\prime}$ represent latent vectors of an accurate autoencoder (that maps related generators $\mathbf{G}$ to close-by vectors), then distance-based losses are natural for comparing $x,x^{\prime}$, as they are sensitive to positional information in latent space. This is not true for, e.g., log-likelihood losses, which are invariant under all measure-preserving transformations. It turns out that if we take log-likelihood-based signaling games loss as formalized, e.g., in (Rita et al., 2022) and we switch to the Wasserstein-type version where the expectation of $|x-x^{\prime}|^{p}$ replaces KL-discrepancy of the distributions of $x,x^{\prime}$ then we directly obtain the basic form of our loss $\mathcal{L}^{int}_{\omega,\phi}$ (see §C for the full derivation). In this sense, the form of the loss (2) is natural for our setting.
4.2 Imitation (Im)
During the Im phase, we apply a teacher-student regime to initialize the next iteration sender $S_{\phi^{t+1}}$. Specifically, frozen $S^{*}_{\phi^{t}}$ and $R^{*}_{\omega^{t}}$ generate ground truth predictions, which $S_{\phi^{t+1}}$ is trained to imitate, in a supervised setting.
Pooling (or mode collapse in the field of generative models) is a common problem in LSG in general and corresponds to the breaking of UE in optimization. As already mentioned in §3.2, this is exacerbated in pretrained generation encoders that map to a continuous space with weak guarantees of separation of the encodings. If there is not enough incentive for the sender to generate separate messages for similar states, then pooling is further favored by the compression pressure applied by our interaction scheme Int, in cases in which a strictly simpler encoding protocol can achieve marginally worse expected reconstruction.
UE requires us to enforce LSG separating equilibria, i.e., encodings that unambiguously represent every possible combination of generating factors. To push the iteration dynamics towards such equilibria, we employ the following three strategies that favor message diversity in the imitated protocol message space:
  1. 1.
    State space imitation: As opposed to other IL implementations (Ren et al., 2020; Li & Bowling, 2019a), the imitation accuracy is not tested in the message space, but rather in the state space, using the receiver network from the interaction phase. This allows more degrees of freedom for the imitation protocol.
  2. 2.
    Final state reconstruction: During the Im phase we use the following reconstruction loss:
    $\mathcal{L}^{rec}_{\phi^{t+1}}:=\mathbb{E}_{x\sim\mathcal{X}}\left[d_{\mathcal{X}}(x^{\star}_{t},x^{\prime}_{t+1})\right],$ (3)
[…]
    Differing from the Int phase, $\mathcal{L}^{rec}_{\phi^{t+1}}$ only includes the final predicted state, associated to the full message of length $C$. Along with the previous point, this should allow the student sender $S_{\phi^{t+1}}$ to more easily find different strategies with the same end result, ideally breaking pooling states encountered during the Int phase.
  3. 3.
    Pairwise distance maximization (PDM): To further favor diversity in the new protocol and encourage exploration, the student maximizes the Hamming distance between all messages in a given batch, pushing the sender to find the most ”diverse” protocol which behaves similarly to the teacher protocol, modulated by a hyperparameter $\beta$. With this, the Im loss becomes:
    $\mathcal{L}^{rec}_{\phi^{t+1}}+\beta\frac{1}{N_{batch}^{2}}\sum_{i,j}D[i,j],$ (4)
    $\text{where }D[i,j]:=d_{H}(S_{\phi^{t+1}}(x_{i}),S_{\phi^{t+1}}(x_{j})).$
    In §B.4 we prove that this simple regularization still gives a lower bound for entropy maximization (Williams & Peng, 1991), and an upper bound on the contrastive learning loss NT-Xtent (Chen et al., 2020).
5 Experiments
5.1 Datasets
We tested our framework using the composition split (Schott et al., 2022) of the Shapes3D (Matthey et al., 2017) dataset, which consists of colored images of 3D geometric shapes, with 6 underlying generating factors $G$: floor hue, wall hue, object hue, shape, scale and orientation. The total amount of atribute-value combinations is $480,000$.
The compositional split ensures that every possible value for each atribute is seen in the training set, yet most combinations of atribute values are reserved for the test set (See §A for a proof that this should hold with high probability).
5.2 Architecture
For the sender and receiver networks, we used the EGG Python library (Kharitonov et al., 2021). Both are single layer LSTMs (Hochreiter & Schmidhuber, 1997b) with an embedding size of 64 and hidden size of 256, followed by a two layer MLP with ReLU activation. We defined a vocabulary size of $|V|=25$, and a fixed channel length of $C=10$. For the channel we used a Gumbel-Softmax bottleneck (Jang et al., 2017) without straight-through estimation.
As a vision backbone, we pretrained a convolutional VAE for 2000 epochs with the Adam optimizer and a learning rate of $10^{-4}$. The encoder uses 4 convolutional layers with kernels in increasing powers of 2, from 32 to 256, with ReLU activation, and a final fully connected layer that outputs the predicted mean and logvariance vectors of dimension 256. The decoder is a mirrored version of the encoder with transposed convolutions, ReLU activation in the hidden layers, and sigmoid activation in the output layer.
Rather than training on the images themselves, we used the pretrained backbone to extract the predicted distributions for the latent vectors, and used these as our states. Since the states are multivariate Gaussians and not deterministic vectors, we take the reconstruction loss over the defined distributions. As mentioned in Rmk. 4.1, taking a latent space distance such as Wasserstein distance ($W_{2}$ in our case) is the most natural approach.
5.3 Training
The interaction-imitation games were run for 500 iterations, alternating one full epoch in each phase per iteration. We found that Int required smaller batch sizes to converge, wheras the Im phase PDM becomes a better entropy approximation at higher batch sizes, however scaling quadratically in computation. Therefore, we set the Int and Im batch sizes to 256 and 512, respectively.
To implement Im we defined a receiver $R_{\omega}$ and two senders $S_{\phi_{1}}$ and $S_{\phi_{2}}$ with their own optimizers. After participating in an Int iteration, $\phi^{1}$ and $\omega$ are frozen and used to train $S_{\phi_{2}}$ as described in §4.2. After Im, the weights $\phi_{2}$ are copied to $\phi_{1}$, and $\phi_{2}$ is reset. Importantly, the optimizers for $\phi_{1}$ and $\phi_{2}$ are never reset nor copied between them.
We compare our model with the following training schemes:
  • Interaction phase only (IPO): Training is performed without resetting the sender between iterations.
  • Interaction phase reset (IPR): After the Int phase, the sender is fully reset and not retrained with Im.
  • IL complete imitation (ILCI): Trained with IL framework using the full predicted sequence in the Im phase.
  • IL final state imitation (ILFI): Trained with IL framework using only the final state of the predicted sequence in the Im phase.
5.4 Metrics
Following (Chaabouni et al., 2022), we utilize the Topographic Similarity (TopSim) (Brighton & Kirby, 2006) between the emergent language and the ground truth generating factors, using Hamming distance for both. Furthermore, we measure the mean Hamming distance over all messages to detect pooling.
To quantify EE, we define the useful length $\ell_{\epsilon}(x)$ as the minimum message length for a given state $x$ and protocol encoded by $S_{\phi},R_{\omega}$, such that the additional accuracy gained by extending the message length by 1 is less than some small $\epsilon$, formally:
${\ell}_{\epsilon}(x):=\min\{i:d_{\mathcal{X}}(R_{\omega}(S_{\phi}(x)_{[i]}),R_{\omega}(S_{\phi}(x)_{[i+1]}))\leq\epsilon\},$
in which we restrict $i$ to $1,\dots,C$ and we take $\ell_{\epsilon}(x)=C$ if the minimum is not achieved in this range. Finally, we use the perceptual distance (Zhang et al., 2018) between generated images and the $W_{2}$ distance on the final predicted vector to measure reconstruction quality.
Figure 3: Comparison of validation reconstruction errors and evolution of useful length $\ell$ during training (averaged over 10 runs).
6 Results
Mode TopSim@200$\uparrow$ Percp. loss@200$\downarrow$
IPO 0.0182 0.4608
IPR 0.2385 0.2183
ILCI 0.2208 0.2126
ILFI 0.2339 0.1903
ILCI+PDM 0.2486 0.1944
ILFI+PDM (ours) 0.2665 0.1900
Table 1: Results from evaluating all tested methods with $\beta=0.5$ and $\lambda=0.95$ over the test set, averaged over 10 runs. To take into consideration that all methods had worsening results given enough iterations, we decided to take these measurements at epoch 200, where most models achieved their best performance.
6.1 Reconstruction performance
As evidenced in Fig. 3, our method achieves the lowest reconstruction error on the test set during early training, and becomes marginally suboptimal with respect to ILFI at the later stages of training. This behaviour is similar to the explore-exploit tradeoff of entropy maximization (Williams & Peng, 1991) in reinforcement learning.
Our method, however, achieves comparatively poor results within the training set, observed in Fig. 3 a), having a comparable full message-loss to IPR. Notably, tested methods without PDM achieve much better results during training, and slightly worse results in test. This behavior confirms our hypotheses about the need for additional inductive biases to enforce generalization in a compositional setting (see §B).
6.2 Impact of IL
As observed in Table 1 and Fig. 3, IL is required for convergence in this setting, as IPO resulted in near absolute pooling. This fact is evidenced by the useful length $\ell$ approaching zero, meaning most states are encoded into the same message.
Surprisingly, IPR consistently converged, which suggests that pooling equilibria are less stable to perturbations of the protocol, even when they are fully random. However, the lack of an Im phase results in lower quality image reconstructions and low TopSim similarity of the learned message space, as seen in Table 1.
6.3 ILCI vs ILFI
As observed in Table 1, ILFI obtains noticeably better results thtan ILCI, as measured by both reconstruction and TopSim. This result confirms our intuition of the importance of increasing the degrees of freedom order to maximize message expressiveness (see §B).
Fig. 3 shows that ILCI and ILFI have a much higher difference in reconstruction accuracy on the test set when not using DPM, behaving similarly on the train set regardless of regularization. This result suggests that utilizing ILFI has a higher impact on generalization than DPM.
6.4 Compression vs generalization
All tested methods converged to a smaller value of $\ell$ during training, with non PDM imitation methods closely approaching the theoretical minimum length $\lceil\log_{|V|}|\mathcal{D}|\rceil=5$ required to unambiguously encode every state. PDM appears to be damping this compression, avoiding full optimization of $\mathbb{E}[\ell]$. We claim this is required for NHE (see §B.3).
By tuning the parameters $\lambda$ and $\beta$, we were able to find a balance between compositionality and compression (See §2). For $\beta$ larger than $1$, the protocol does not converge efficiently, behaving in a similar manner to IPR. With $\lambda=0.95$ and $\beta=0.5$, we obtained the highest TopSim and the smallest perceptual loss on the training set among all tested methods (see Table 1).
7 Limitations and future work
As this work is the first proposal of an LSG + IL framework for generation tasks, several directions for future work emerged, and are outside the scope of this work:
  1. 1.
    For ease of rigorous evaluation, we performed a large ablation study on a toy dataset. In future work it is interesting to consider real-world data in the same framework. This will require an enriched (and plausibly valuable in itself) discussion of new ad-hoc evaluation metrics.
  2. 2.
    We considered a product-type compositional hypothesis, i.e., generating factors are assumed to be the outcome of independent components $\mathbf{G}=[G_{1},\dots,G_{n}]$. It is interesting to extend the theoretical framework (§A and 3.1) to more general compositional structures.
  3. 3.
    We assumed deterministically generated data as an injective function of $\mathcal{G}$. A clear avenue for future study lies in using a more general $\mathsf{GenX}$ as a randomized function with noise input as well as $\mathcal{G}$.
  4. 4.
    Generative tasks permit evaluations of novel compositional properties. We qualitatively explored the notable case message interpolation and found interpretable results (see §F.1 and §F.2). Exploring making quantitative such evaluations, seems a valuable avenue of future research in compositional generative machine learning.
8 Conclusion
In this work we presented a novel self-supervised framework for enforcing compositionality in generative models, applicable to probabilistic generative models such as VAEs.
By leveraging emergent language models, we created a tunable comunication bottleneck between collaborative agents, with ad-hoc biases towards compressibility and expressibility. Through this bottleneck, the latent vectors of probabilistic models could be compressed to a message space that shows good topological similarity with the hidden underlying generating space $\mathcal{G}$ of the dataset, indicating that the produced encodings inherited an approximate compositional structure.
We provided rigorous justifications and empirical evidence, in the form of a detailed ablation study, for the impact of our framework choices on the emergent properties of the comunication protocols.
Appendix A Recovering compositional structure from partially observed data
For the sake of clarity of treatment, we assume here that $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each factor $G_{i}$ has the same number $N$ of possible values, and as a consequence $\mathcal{G}=[1:N]^{n}$. Our results will extend directly to the general $\mathcal{G}$ case as shown in Rmk. A.4, but that statement seems too tedious for the reader to parse.
We work under the simplified deterministic generation hypotheses, in which our dataset is $\mathcal{D}=\mathsf{GenX}(\mathcal{G})\subset\mathcal{X}$ in which $\mathsf{GenX}$ is a deterministic injective function.
Our main result is the following:
Theorem A.1.
Assume that $\mathcal{G}=[1:N]^{n}$ and that $\mathsf{GenX}:\mathcal{G}\to\mathcal{X}$ is an injective function, and that $\mathcal{D}_{train}\subseteq\mathcal{D}$ has cardinality $|\mathcal{D}_{train}|=p|\mathcal{D}|$ for some $p\in(0,1)$ such that $pN^{n}$ is an integer.
We assume that the compositional structure of $\mathcal{G}$ is such that the following holds:
  1. (A)
    For some $k<n$, if we observe a set of data whose generating factor combinations feature all possible combinations of generating factor values $(G_{i_{1}},\dots,G_{i_{k}})$ for all choices of indices $1\leq i_{1}<\cdots<i_{k}\leq n$ then this is sufficient to reconstruct $\mathcal{G}$ to good accuracy.
Then as $|\mathcal{G}|\to\infty$ the probability that $\mathcal{D}_{train}$ is sufficient to reconstruct $\mathcal{G}$ tends to $1$.
Here (A) makes rigorous the common assumption that having a training set that features examples for diverse enough combinations of factors is sufficient for reconstructing all combinatorial generators $\mathbf{G}$.
Note that in this work we do not discuss
  1. 1.
    what features of the pretrained encodings $\mathcal{G}\to\mathcal{D}$ and what requirements on the compositional structure of $\mathcal{G}$ would allow to actually guarantee this assumption,
  2. 2.
    alternatives for a rigorous definition of what is meant by ”to good accuracy” in the statement of assumption (A).
Once the above points are settled and defined, we can replace (A) by an explicit requirement. Since however these points seem like an ambitious whole research line and are far beyond the scope of this work, we leave the definition of ”to good accuracy” and the proof of reconstruction to future works.
Thus, introducing (A) allows us to prove fully rigorous statements such as Thm. A.3, in absence of a full theory of compositional reconstructon.
Nevertheless, we make the following remarks:
  • There is a natural trade-off between assumptions on how big $\mathcal{D}_{train}$ should be, vs. how complicated the compositional reconstruction should be allowed to be for reconstruction to be feasible: the richer set of such combinations we require $\mathcal{D}_{train}$ to contain (thus making more restrictive assumptions on $\mathcal{D}_{train}$), the lower is the requirement for our encoding framework and for the compositional structure for (A) to hold.
  • It seems natural to conjecture that if assumption (A) does not hold for large value of $k$, then there is no hope for the compositional reconstruction to be achievable. Thus lowering the requirement on ”good enough accuracy” to make (A) hold seems the only avenue of future research.
Theorem A.1 is a consequence of below Theorem A.3 that proves that assumption (A) holds with probability tending to $1$ as $|\mathcal{G}|\to\infty$ and $|\mathcal{D}_{train}|/|\mathcal{G}|=p$ is bounded away from zero. What Theorem A.3 says is that for large sets of generators, it will become overwhelmingly unlikely that $\mathcal{D}_{train}$ does not contain enough data to reconstruct the generators.
A.1 Observing $\mathcal{D}_{train}$ allows to observe finite statistics of the factors with high probability
Our first result for this section is that when $\mathcal{D}_{train}\subseteq\mathcal{D}$ is chosen uniformly at random amongst sets of size $p|\mathcal{D}|$, for $|\mathcal{D}|=|\mathcal{G}|$ large enough, we find that with overwhelming probability all possible values of each of the factors $G_{1},\dots,G_{n}$ are achieved by some elements of $\mathcal{D}_{train}$. We state the result in Proposition A.2, whose proof in our view contains the main ideas useful also for the full result of Theorem A.3, which can be considered our main technical result for this section.
The first result is as follows:
Proposition A.2.
Assume that for $n,N\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N\}=[1:N]^{n}$ and that $\mathsf{GenX}:\mathcal{G}\to\mathcal{X}$ is an injective function with image $\mathcal{D}$ as above. Denote by $\mathcal{G}_{train}:=\mathsf{GenX}^{-1}(\mathcal{D}_{train})$ the generating factors related to the training set, and assume that $\mathcal{D}_{train}\subset\mathcal{D}$ is chosen uniformly at random amongst subsets of cardinality $p|\mathcal{D}|$, for some $p\in(0,1)$ such that $pN^{n}$ is an integer. Then
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}\leq nN\left(1-\frac{1}{N}\right)^{pN^{n}},$ (5)
in particular at fixed $p$, the above event has probability tending to zero as $N^{n}\to\infty$.
Thus, in an equivalent formulation of the last part of the above proposition, we have that in the limit of large $|\mathcal{G}|=N^{n}$, the probability that each generating factor $G_{i}$ takes all of its possible values on some generated element of $\mathcal{G}_{train}$ tends to $1$.
Proof.
The possible choices of $\mathcal{G}_{train}\subseteq\mathcal{G}$ of cardinality $p|\mathcal{G}|=pN^{n}$ are $\tfrac{(N^{n})!}{(pN^{n})!((1-p)N^{n})!}$. Then, calling $A_{ij}$ the set of elements $\mathbf{G}\in\mathcal{G}_{train}$ such that $G_{i}\neq j$, we find that $|A_{ij}|=N^{n-1}(N-1)=N^{n}\tfrac{N-1}{N}$, so that the number of possible choices of $\mathcal{G}_{train}\subseteq A_{ij}$ of cardinality $pN^{n}$ is given by $\tfrac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$. Now we bound the probability of the event on the left hand side in (5) via a union bound over the $Nn$ possible choices of $A_{ij}$ for $1\leq i\leq n,1\leq j\leq N$, so that
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}$$\leq$ $nN\frac{(pN^{n})!((1-p)N^{n})!}{(N^{n})!}\frac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$
$=$ $nN\frac{N^{n}\frac{N-1}{N}}{N^{n}}\frac{N^{n}\frac{N-1}{N}-1}{N^{n}-1}\cdots\frac{N^{n}\frac{N-1}{N}-pN^{n}+1}{N^{n}-pN^{n}+1}$
$\leq$ $nN\left(1-\frac{1}{N}\right)^{pN^{n}},$
in which in the first line we used a union bound together with the computations of combinatorial coefficients from above, in the second line we performed a simplification and reordering, and in the third line we used the fact that $1-1/N=(N^{n}(N-1)/N)/N^{n}>(N^{n}(N-1)/N-i)/(N^{n}-i)$ for all $11\leq i<pN^{n}-$ 1.
Finally, considering the limit $N^{n}\to\infty$ of the upper bound in (5), we have two cases: (1) if $N\geq 2$ stays bounded and $n\to\infty$, then we can bound $nN(1-1/N)^{pN^{n}}\leq C_{1}nC_{2}^{C_{3}^{n}}$ with constants $C_{2}\in(0,1)$ and $C_{3}>1$, and by taking logarithms we check that the quantity tends to zero; (2) if $N\to\infty$ and $n\geq 2$, and in this case $(1-1/N)^{pN^{n}}\sim e^{-pN^{n-1}}$ and again by taking logarithms we find that the quantity tends to zero. ∎
In fact Prop. A.2 is gives us guarantees that it will be possible to observe all options for each $G_{i}$ in the training set, however this may not be sufficient for the main aim of recovering the compositional structure of $\mathcal{G}$. However we next prove, in an extension of Prop. A.2, that actually we will also observe with high probability any possible combination of finitely many factors:
Theorem A.3.
Under the same hypotheses on $\mathcal{G},\mathsf{GenX},\mathcal{G}_{train},p$ as in Prop. A.2, for any $k\leq n$ we have the following.
$\mathbb{P}\{(\exists i_{1}<\dots<i_{k}\leq n)(\exists j_{1},\dots,j_{k}\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\}$
$\leq\frac{n!}{k!(n-k)!}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}},$ (6)
which for fixed $p,k$ tends to zero as $N^{n}\to\infty$.
Proof.
As mentioned before, roughly the overall proof strategy is the same as for Prop. A.2. We introduce the sets $A^{j_{1},\dots,j_{k}}_{i_{1},\dots,i_{k}}=A^{\mathbf{j}}_{\mathbf{i}}$ and observe that the number of choices of $i_{1}<\cdots<i_{k}\leq n$ is $\tfrac{n!}{k!(n-k)!}$ while the number of choices of $j_{1},\dots,j_{k}\leq N$ is $N^{k}$. Thus the number of sets of the form $A^{\mathbf{j}}_{\mathbf{i}}$ is $N_{A}:=\tfrac{N^{k}n!}{k!(n-k)!}$. Furthermore, each set $A^{\mathbf{j}}_{\mathbf{i}}$ only imposes the constraint of entries of indices $i_{1},\dots,i_{k}$ from a $\mathbb{G}$ to avoid one combination of values $\vec{j}$, and thus it has cardinality $|A^{\mathbf{j}}_{\mathbf{i}}|=N^{n-k}(N^{k}-1)$. Then the number of choices of $\mathcal{G}_{train}\subseteq A^{\mathbf{j}}_{\mathbf{i}}$ of cardinality $pN^{n}$ is just the combinatorial coefficient $\tfrac{(N^{n-k}(N^{k}-1))!}{(pN^{n})!(N^{n-k}(N^{k}-1-pN^{k}))!}$ and by union bound we then find by the same reasoning as in Prop. A.2,
[…]
$=$ $N_{A}\frac{N^{n}-N^{n-k}}{N^{n}}\frac{N^{n}-N^{n-k}-1}{N^{n}-1}\cdots\frac{N^{n}-N^{n-k}-pN^{n}+1}{N^{n}-pN^{n}+1}$
$\leq$ $N_{A}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}=\frac{n!}{k!(n-k)!}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}.$
This proves (A.3). In order to show that the bound tends to zero as $N^{n}\to\infty$ at fixed $p>0,k\leq n$, we first note that $\tfrac{n!}{k!(n-k)!}\leq n^{k}/k!$ and thus (as $k$ is assumed fixed) it suffices to show that
$n^{k}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}\to 0\quad\text{as}\quad N^{n}\to\infty.$
Here again, we can proceed with the discussion of the two cases ($N$ bounded and $n\to\infty$, or $N\to\infty$) exactly as in the end of the proof of Prop. A.2, to conclude. ∎
Remark A.4 (Extension of Prop. A.2 and Theorems A.3 and A.1 to general $\mathcal{G}$).
If, for a choice of $n\geq 2$ and $N_{1},\dots,N_{n}\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N_{i}\}$, then $|\mathcal{G}|=N_{1}\cdots N_{n}$ and by carefully following the same steps as in the proof of Prop. A.2 we find
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N_{i})(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}\leq\sum_{i=1}^{n}N_{i}\left(1-\frac{1}{N}_{i}\right)^{pN_{1}\dots N_{n}},$ (7)
Again, this tends to zero for fixed $p$ as $|\mathcal{G}|=N_{1}\dots N_{n}\to\infty$: to prove it, we can proceed as follows. In a first case, if $n>2$ stays bounded, then we can discuss separately for each $1\leq i\leq n$ whether $N_{i}\to\infty$ or whether it stays bounded, as done in the proof of Prop. A.2, and in either case the corresponding summand on the left of (5) tends to zero, as in case (2) from the proof of Prop. A.2. Otherwise, $n\to\infty$ and we can proceed analogously to case (1) from the end of proof of Prop. A.2.
For the extension of Thm. A.3 to general $\mathcal{G}$ we get the bound
$\mathbb{P}\{(\exists i_{1}<\dots<i_{k}\leq n)(\exists j_{1},\dots,j_{k}\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\}$
$\leq\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}N_{i_{1}}\cdots N_{i_{k}}\left(1-\frac{1}{N_{i_{1}}\cdots N_{i_{k}}}\right)^{pN_{1}\cdots N_{n}},$ (8)
and the strategy for its proof and for the proof that this goes to zero if $k,p$ fixed and $|\mathcal{G}|=N_{1}\cdots N_{n}\to\infty$ are a direct extension of the above.
Theorem A.1 is a direct consequence of assumption (A) and Theorem A.3, and thus it also directly extends to the general case.
Appendix B Regularization to counteract Holistic Encodings
In this section, we include rigorous results useful to justify the regularization of our Im phase, aimed at balancing NHE and EE.
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. (Kirby & Hurford, 2002; Kirby et al., 2015)). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
Our aim in this section is threefold:
  1. 1.
    In §B.1 we set rigorous definitions of Holistic Encodings within the setting of §2.3 and §3.1.
  2. 2.
    In §B.2 we check that NHE is required for lowest Kolmogorov Complexity encodings.
  3. 3.
    We verify that optimizing for just UE and EE on our training set, will produce Holistic Encodings, and thus is not satisfactory for reconstructing $\mathcal{G}$ (see §B.3);
  4. 4.
    We give a few possible options on how to enrich the loss functions in our IL setup, aimed at establishing a better balance between EE and NHE (see §B.4).
B.1 Definitions of Holistic and Compositional encodings, and role of the NHE condition
In the classical IL setup (Kirby & Hurford, 2002; Kirby et al., 2015), a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but compl forgetting the compositional structure of the ”language” to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
Definition B.1.
In the setting from §2.3 and §3.1, sender $S:\mathcal{X}\to\mathcal{M}$ determines a holistic encoding if it satisfies UE but the messages associated to data generated as $\mathsf{GenX}(\mathbf{G})$ do not have a compositional structure in terms of generating factors $\mathbf{G}$, as specified by Definition B.2 below.
Note that, as in (Kirby & Hurford, 2002; Kirby et al., 2015), Definition B.1 of holistic encodings refers to ”respecting compositional structure”, which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
Definition B.2.
In the setting of §2.3 and §3.1, assume that generating factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ have $G_{i}\in[1:N_{i}]$ for some $N_{1},\dots,N_{n}\geq 2$. Then sender $S:\mathcal{X}\to\mathcal{M}=V^{C}$ will be defined to respect the structure of generating factors $\mathbf{G}$ if for each $1\leq i\leq n$ there exist an injective mapping $E_{i}:[1:N_{i}]\to V^{C_{i}}$, for which
$\forall\mathbf{G}\in\mathcal{G},\ S(\mathsf{GenX}(\mathbf{G}))=E_{1}(G_{1})\cdots E_{n}(G_{n})\quad\text{ and }\quad\sum_{i=1}^{n}C_{i}=C.$
As indicated in §3.4, NHE condition on the sender-receiver protocol just requires that the protocol be not holistic, and thus it is fully specified via definitions B.1 and B.2. Furthermore, Definition B.2 fully specifies in which way the map $\mathcal{G}\to\mathcal{M}$ successfully approximates an isomorphic reconstruction of $\mathcal{G}$, if it satisfies conditions NHE and UE.
B.2 Encodings optimizing for only UE and EE have Kolmogorov Complexity much higher than compositional ones
By the classical learning theory principle, if our sender generates encodings with lower complexity, these should be easier to learn by the receiver, in the sense that the receiver will tend to generalize with more ease form good accuracy on the training set to similar accuracy on the test set.
As a manageable measure of complexity we will use Kolmogorov Complexity (KC), and we verify that in our setting, encodings satisfying UE and EE will necessarily have higher KC than compositional ones. In particular, the receiver will have more trouble reconstructing from such encodings, justifying our push to enforce the NHE condition.
Encodings of $\mathcal{D}_{train}$ satisfying UE and EE include all injective maps
$\mathsf{enc}:\mathcal{D}_{train}\to V^{C_{\star}},\quad C_{\star}=\lceil\log_{|V|}|\mathcal{D}_{train}|\rceil,$ (9)
in which $\lceil a\rceil$ denotes the smallest integer larger than real number $a$. For a random map $\mathsf{enc}$ as above, the encoding to be Kolmogorov-irreducible, so that the expected KC of a random $\mathsf{enc}$ would be up to constant factor the one corresponding to explicitly enumerating the $|\mathcal{D}_{train}|$ strings of length $C_{\star}$ in alphabet $V$ that describe the encodings of each element of the training set. Each such string thus requires $\approx C_{\star}\log_{2}|V|$ bits. Thus if $U$ is the uniform distribution over injective maps (9) then we have
$\mathbb{E}_{\mathsf{enc}\sim U}[KC(\mathsf{enc})]\approx|\mathcal{D}_{train}|C_{\star}\log_{2}|V|\approx|\mathcal{D}_{train}|\log_{2}|\mathcal{D}_{train}|.$ (10)
On the other hand, to specify a compositional encoding $\mathsf{comp}$ as in Definition B.2 requires only assigning a string of length $C_{i}$ in alphabet $V$ for every generating factor $G_{i}$, where we can take $C_{i}=\lceil\log_{|V|}N_{i}\rceil$. Thus such string requires $\approx\log_{2}N_{i}$ bits, and the whole encoding has KC given by
$KC(\mathsf{comp})\approx\sum_{i=1}^{n}\log_{2}N_{i}=\log_{2}|\mathcal{D}_{train}|.$ (11)
If we assume that $|\mathcal{D}_{train}|/|\mathcal{D}|=p\in(0,1)$ and we take $|\mathcal{D}|\to\infty$, then we get
[…]
The computation (12) shows that
Proposition B.3.
Under the same hypotheses on $\mathcal{G}$ as in the previous section, consider the limit $|\mathcal{D}|\to\infty$ and assume that $p=|\mathcal{D}_{train}|/|\mathcal{D}|$ stays bounded away from zero in the limit. Then the expected Kolmogorov complexity of encodings satisfying UE and EE becomes overwhelmingly larger than the one of compositional encodings:
$\frac{\mathbb{E}_{\mathsf{enc}\sim U}[KC(\mathsf{enc})]}{KC(\mathsf{comp})}\sim p|\mathcal{D}|\to\infty.$
B.3 Encodings that optimize efficiency of encoding EE under UE over $\mathcal{D}_{train}$ are holistic
Example 3.1 from §3.4 may be sufficient to convince the reader intuitively of this phenomenon, however here we develop the idea of that example in higher formal detail.
First, we clarify the definitions:
  • The requirement of ”UE over $\mathcal{D}_{train}$” means that the encodings are injective over $\mathcal{D}_{train}$.
  • The requirement of ”EE over $\mathcal{D}_{train}$” means that the encoding maps $\mathcal{D}_{train}$ into $V^{C_{\star}}$ for the minimum value of $C_{\star}$.
Similarly to previous subsection (see (9)), keeping in mind that $|\mathcal{D}_{train}|=p|\mathcal{D}|=p|\mathcal{G}|$, with fixed $p\in(0,1)$, we have
$C_{\star}=\lceil\log_{|V|}|\mathcal{D}_{train}|\rceil=\lceil\log_{|V|}p|\mathcal{G}|\rceil\leq\lceil\log_{|V|}|\mathcal{G}|\rceil-\lfloor|\log_{|V|}p|\rfloor.$ (13)
For compositional encodings of $\mathbf{G}=[G_{1},\dots,G_{n}]$ the encoding will require to set apart separate message substrings of length $C_{i}=\lceil\log_{|V|}N_{i}\rceil$ for factor $G_{i}$ (see Def. B.2), thus the required length for the encoding will be
$C_{0}:=\sum_{i=1}^{n}\lceil\log_{|V|}N_{i}\rceil\stackrel{{\scriptstyle(\star)}}{{\geq}}\left\lceil\sum_{i=1}^{n}\log_{|V|}N_{i}\right\rceil=\lceil\log_{|V|}|\mathcal{G}|\rceil\stackrel{{\scriptstyle(\star\star)}}{{\geq}}C_{\star}+\lfloor|\log_{|V|}p|\rfloor,$ (14)
[…]
We now discuss when the inequality signs in (14) are sharp or not, since sharp inequality $C_{0}>C_{\star}$ implies our claim that optimizing for EE under UE implies non-compositional (i.e., holistic) encodings:
  1. 1.
    If we use relatively small $|V|\lesssim 10$ (assuming commonly used values $p\lesssim 0.1$), we will have
    $|V|\leq\frac{1}{p},$ (15)
[…]
  2. 2.
    If the gaps between $\log_{|V|}N_{i}$ and the lowest integer larger or equal to it, sum to a value of at least $1$, i.e.,
    $\sum_{i=1}^{n}\left(\lceil\log_{|V|}N_{i}\rceil-\log_{|V|}N_{i}\right)\geq 1,$ (16)
    then ()($\star$)( ⋆ ) becomes sharp, again guaranteeing $C_{0}>C_{\star}$ as desired. This could be probable to happen in our setting, and becomes more likely for large $|V|>1/p$ for which (15) fails, given that the $N_{i}$ are supposed to be unknown, so it is likely that for large $|V|>1/p$ for a few of the $N_{i}$ the gaps in (16) are nontrivial. For example, (16) holds true for $|V|>1/p\geq 10$ if, say, at least two of the $G_{i}$ are binary factors, so that $N_{i}=2$.
In summary we have the following:
Proposition B.4.
Under the same running assumptions over $\mathcal{G}$, $p$, we have that optimizing only for efficiency of encodings EE under injectivity constraints UE over $\mathcal{D}_{train}$ is guaranteed to produce holistic (i.e., non-compositional) encodings if one or both of the conditions (15) and (16) holds.
B.4 Possible choices of encoded message regularization
As seen above, if we just optimize for EE and UE then we are likely to get holistic encodings (Prop. B.4), which will then make it hard for the receiver to generalize the decoding strategy outside the training set due to having overwhelmingly higher expected complexity (Prop. B.3).
This proves that it is important to incentivize condition NHE, i.e., to push encoding strategies away from overly efficient holistic encodings. In this section we discuss a few approaches to do this in practice, explaining our choice of NHE regularization.
  1. 1.
    Entropy maximization. Introduced by (Williams & Peng, 1991) for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
    $h(\xi,W,x)=-\ln Pr(y=\xi|W,x),$ (17)
    such that if we take the expectation over $\xi$ we get:
    $\mathbb{E}[h(\xi,W,x)|W,x]=-\sum_{\xi}Pr(y=\xi|W,x)\ln Pr(y=\xi|W,x),$ (18)
    which is the entropy of the network. Therefore $h$ is an unbiased estimator of the entropy. They also note that, if $w_{i}$ and $x_{i}$ with $i\in U$ represent the weights and pattern of single neurons in a feed forward network, $\xi_{i}$ the $i_{th}$ position in a $n_{u}$-tuple $\xi$, and let $g_{i}$ be the probability density describing $y_{i}$, i.e:
    $g_{i}(\xi,w^{i},x^{i})=Pr(y_{i}=\xi|w^{i},x^{i}),$ (19)
    then, for any feed forward network,
    $Pr(y=\xi|W,x)=\prod_{i\in U}g_{i}(\xi_{i},w^{i},x^{i}),$ (20)
    and therefore
    $-\ln Pr(y=\xi|W,x)=-\sum_{i\in U}g_{i}(\xi_{i},w^{i},x^{i})=h(y,W,X).$ (21)
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in (Ren et al., 2020).
    Naive adaptation to our setting is not practical. We take $x\in\mathcal{X}$ and $\xi\in\mathcal{M}=V^{C}$ and recall that in our notation $m_{t}$ indicates the $t$-th token of message $m$, and $m_{[t]}$ is the notation for the initial segment $m_{1},m_{2},\dots,m_{t}$, so that in particular $m_{[C]}=m$.
    Importantly, in our setting, to avoid high variance in estimation due to the large message space (see §C), we approximate the probability distribution of the RNN-produced message token at position $t\leq C$, denoted here $p(m_{t}|x,m_{[t-1]})$, as a $\delta$ distribution over the vocabulary $V$ by using the Gumbel-softmax trick:
    $(\exists v\in V),\quad Pr(m_{t}\mid x,W,m_{[t-1]})\approx\begin{dcases}1,&\text{ if }m_{t}=v\\ 0,&\text{ if }m_{t}\in V\setminus\{v\},\end{dcases}$ (22)
    and therefore:
    $Pr(m=\xi|W,x)=\prod_{t=1}^{C}Pr(m_{t}=\xi_{t}|W,x,m_{[t-1]}).$ (23)
    Note that due to (22), for all but a single one of the $|V|^{C}$ possible values of $\xi$ the quantity (23) is $\approx 0$, and thus $h$ becomes a poor estimator for the entropy of $p(m_{t}|x,m_{[t-1]})$, as we get
    $h(m,W,x)=-\ln\sum_{t=1}^{C}Pr(m_{t}=\xi_{t}|x,W,m_{[t-1]})\approx\infty.$ (24)
    In particular, it becomes clear that optimizing this $h$ becomes unfeasable with gradient methods.
    Adaptation to our setting: a more promising strategy. Our model can be seen as a different approximation of the entropy of the message. Instead of using the policy weights as probability values, we use the law of large numbers to estimate the probability of each message. For a batch $\mathcal{B}\subset\mathcal{X}$ consisting of $N_{batch}$ training examples and for given sender encoding protocol $S_{\phi}$:
    For a sufficiently large $\text{For a sufficiently large $N_{batch}$:}\quad\Pr(m=\xi\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)=\xi\}|}{N_{batch}}\$ (25)
    Yet again, due to the exponential size of the message space, the entropy taken by this probability is still not practical. As an approximation, we will asume independence between symbol positions, allowing to get the stronger approximation:
    $\Pr(m_{t}=v\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)_{t}=v\}|}{N_{batch}},$ (26)
    as well as the additivity of entropy to be applied to (18):
    $H(Pr(m|x,\phi))$ ::=:= $-\sum_{\xi\in V^{C}}Pr(m=\xi\mid x,\phi)\ln Pr(m=\xi\mid x,\phi)$ (27)
    $=$ $-\sum_{t=1}^{C}\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\ln Pr(m_{t}=v\mid x,\phi).$
    Then we can perform the following computations, in which we let $m,m^{\prime}$ be two i.i.d. copies of $m$:
    $H(Pr(m\mid x,\phi))$ ${{=}}$ $-\sum_{t=1}^{C}\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi))\ln(Pr(m_{t}=v\mid x,\phi))$ (28)
    $\geq$ $\sum_{t=1}^{C}\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi)-Pr(m_{t}=v\mid x,\phi)^{2})\quad\text{(because $-x\ln x\geq x-x^{2}$ for $x\leq 1$)}$
    $\stackrel{{\scriptstyle(\star)}}{{=}}$ $\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)-\sum_{v\in V}(Pr(m_{t}=v\mid x,\phi)^{2})\right)$
    $=$ $\sum_{t=1}^{C}\sum_{v\neq v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)=\sum_{t=1}^{C}Pr(m_{t}\neq m_{t}^{\prime})=\mathbb{E}[d_{H}(m,m^{\prime})]$
    (${{\approx}}$ $\mathbb{E}_{x,x^{\prime}\in\mathcal{B}}[d_{H}(S_{\phi}(x),S_{\phi}(x^{\prime})],\quad\quad\text{whith $\mathbb{E}$ taken w.r.t. the uniform measure over }\mathcal{B}.$
    In the above step ()($\star$)( ⋆ ) we used that
    $\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)=1=\left(\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\right)^{2}=\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi).$
    The consequence of bound (28) is that maximizing the pairwise Hamming distance between predicted messages in a batch also maximizes a lower bound to the approximate entropy of the message space, justifying the choice of optimizing this quantity (which is easier to i include in practice) for favoring NHE.
  2. 2.
    Contrastive Learning losses. Contrastive loss functions such as the triplet loss (Schroff et al., 2015) and NT-Xtent (Chen et al., 2020) work by defining positive and negative image pairs (usually through image augmentation), and aiming to maximize the distance between embeddings of positive pairs and minimize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair for each $x$ in a minibatch $N$ by augmenting each image. Each image in the batch therefore has $12N-$ 1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
    $\ell_{i,j}=-\log\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum^{2N}_{k}\exp(sim(z_{i},z_{k})/\tau)}$ (29)
    For simplicity, let $\tau=1$ and thus
    $\ell_{i,j}=\log\sum^{2N}_{k}\exp(sim(z_{i},z_{k}))-sim(z_{i},z_{j})$ (30)
    If we take this loss over an unaugmented batch, i.e the only application of this loss is for $z_{i}$ with itself, then it reduces to:
    $\ell_{i}=\log\sum^{N}_{K}\exp(sim(z_{i},z_{k}))-1$ (31)
    In our representations, the embedding vectors $z$ are one-hot matrices representing a string of characters. For one hot matrices $A_{C\times N}$ and $B_{C\times N}$, the dot product $AB^{T}$ is equivalent to counting the amount of mathcing rows, i.e, the un-normalized Hamming distance, and both matrices have norm $N$:
    $sim(A,B)=A^{T}B/\|A\|\|B\|=\frac{1}{N^{2}}\sum_{i}^{N}\mathds{1}[argmax(A_{i})=argmax(B_{i})]=\frac{C}{N^{2}}(1-d_{H}(A,B))$ (32)
    , in which as before $d_{H}$ is the Hamming distance. We can replace this expresion in the unaugmented NT-Xtent loss:
    $\ell_{i}$ $=$ $\log\sum_{k}^{N}\exp(\frac{C}{N^{2}}(1-d_{H}(z_{i},z_{k}))-1=\log\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))+\frac{C}{N^{2}}-1$
    $\leq$ $\log\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))=\log N+\log\frac{1}{N}\sum_{k}^{N}\exp(-\frac{C}{N^{2}}d_{H}(z_{i},z_{k}))$
    $\stackrel{{\scriptstyle\text{(Jensen ineq.)}}}{{\leq}}$ $\log N-\frac{C}{N^{3}}\sum_{k}^{N}d_{H}(z_{i},z_{k}).$
    As a consequence, if $C\leq N^{2}$ then $\ell_{i}\leq\log N-\frac{1}{N}\sum_{k}^{N}d_{H}(z_{i},z_{k})$. Thus, for each element $z_{i}$ in a batch $N$, the negative Hamming distance between $z_{i}$ and all other elements plus the logarithm of the batch size form an upper bound on the NT-Xtent loss for the same element, and minimizing the negative Hamming distance over all elements (or equivalently maximising the positive distance) also minimizes the NT-Xtent loss.
Appendix C Justification of our imitation step starting from classical probabilistic reconstruction games
We start by recalling classical probabilistic reconstruction games, after which we adapt the formulation to the setting of generative tasks, and we show that with a few justified simplifications we arrive to our formulation from the main paper.
C.1 Classical probabilistic reconstruction games
Recall that in reconstruction games, the two agent types are a teacher and student (or speaker and listener, or sender and receiver), parameterized by respectively $\theta,\phi$. The game consists in the teacher observing a signal $x$ distributed according to a probability distribution $p$, from which it produces a randomized message $m$ whose distribution conditional on $x$ is obtained according to policy $\pi_{\theta}(\cdot|x)$. The student or listener, observes message $m$ and produces a reconstructed signal distributed according to a conditional probability $\rho_{\phi}(\cdot|m)$. The game dynamics can be modelled by stipulating that together, teacher and student minimize the log-likelihood loss function
$\mathcal{L}_{\theta,\phi}=-\mathbb{E}_{x\sim p,m\sim\pi_{\theta}(\cdot|x)}\left[\log\rho_{\phi}(x|m)\right]$ (33)
in which $x$ represents the signal and $m$ the message used to reconstruct it, the speaker produces a reinforcement learning policy $\pi_{\theta}(\cdot|x)$ and the listener produces a reconstruction function $\rho_{\phi}(x|m)$.
C.2 Reconstruction games in generative tasks
While the above formulation (33) is formalized using the log-likelihood between distributions, for our generative task, the reconstruction modelled slightly differently: the task is to optimize, with high probability, the reconstruction error, measured by a distance between the signal $x$ and its reconstruction $x^{\prime}$, denoted $d(x,x^{\prime})$. Here the notation for $x,x^{\prime}$ and $d(x,x^{\prime})$ can be (ab)used to represent two alternative settings: either (a) $x,x^{\prime}$ are vectors in the autoencoder latent space and $d(x,x^{\prime})$ is a discrepancy (e.g. $d(x,x^{\prime})=|x-x^{\prime}|^{p}$) between them or (b) $x,x^{\prime}$ are reconstructed images and $d(x,x^{\prime})$ is a discrepancy (e.g. based on some ad-hoc norm or otherwise) between them. Thus we have the probabilistic error function
$\mathcal{L}_{int}^{\mathbb{P}}(\omega,\varphi):=\mathbb{E}_{x\sim p,m\sim\pi_{\varphi}(\cdot|x),x^{\prime}\sim\rho_{\omega}(\cdot|m)}d(x,x^{\prime}).$ (34)
The loss (34) is akin to a $d$-based Wasserstein distance analogue of (33).
C.3 The problem reduces to the case of a deterministic student
Note that when $d(x,x)=|x-x^{\prime}|^{2}$ for some Euclidean norm (or more generally, when $d$ is a convex function of such norm), then the minimization of (34) over distributions $\rho_{\omega}$ at fixed $m,\varphi$ must be achieved by
$x_{m}^{\varphi}\in\mathrm{argmin}_{x^{\prime}}\mathbb{E}_{x\sim p_{\varphi}(\cdot|m)}d(x,x^{\prime}),$ (35)
where if the set of $x$ is finite then $p_{\varphi}$ is the probability defined by
$p_{\varphi}(x^{\prime}|m)=\frac{p(x^{\prime})\pi_{\varphi}(m|x^{\prime})}{\sum_{y}p(y)\pi_{\varphi}(m|y)},$ (36)
and an analogous expression with integrals replacing the sum holds in general.
Under the above convexity hypotheses, the choice of $x_{m}^{\varphi}$ in (35) is unique, and for $d(x,x^{\prime})=|x-x^{\prime}|^{2}$ it coincides with the barycenter of $p_{\varphi}$. In this case, $\rho_{\omega}(\cdot|m)$ can be taken to be the Dirac distribution concentrated at $x_{m}^{\varphi}$, and optimization of (34) can be achieved by a deterministic reconstruction function $R_{\omega}(m)$, whose optimum value for given $\varphi$ will be $x_{m}^{\varphi}$ from (35). When reducing to optimizing only amongst deterministic $R_{\omega}$, the loss (34) rewrites in the simplified form
$\mathcal{L}_{int}(\omega,\varphi)=\mathbb{E}_{x\sim p,m\sim\pi_{\varphi}(\cdot|x)}d(x,R_{\omega}(m)).$ (37)
C.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in (Havrylov & Titov, 2017) and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE (Williams, 1992), however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick (Jang et al., 2017), in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (37) as
$\mathcal{L}_{int}(\omega,\phi)=\mathbb{E}_{x\sim p}d(x,R_{\omega}(S_{\phi}(x))),$ (38)
which for $d(x,x^{\prime})=|x-x^{\prime}|$ is very similar to our actual interaction loss (2):
$\mathcal{L}_{int}(\omega,\phi)=\mathbb{E}_{x\in\mathcal{X}}\left[|x-R_{\omega}(S_{\phi}(x))|\right].$ (39)
C.5 Further ad-hoc improvements to the loss
The actual further improvements of (2) compared to (38) are twofold, and for them our motivation is to bias the model towards higher compositional structure of our message encoding:
  1. 1.
    We use an LSTM architecture for $S_{\phi}$, in which the later words of message $m$ have positional dependency on earlier words. Furthermore, as information is fed back to the model in successive LSTM steps, by the signal processing mutual information inequality, as observed within the Information Bottleneck Principle (Tishby et al., 2000; Shwartz-Ziv & Tishby, 2017) earlier words of a message will have higher mutual information with the input than later words. This signal processing mutual information inequality is a first way in which we bias the model towards condensing higher information content in the first words of the messages.
  2. 2.
    As the space of possible messages is huge, and keeping in mind the simplicity bias of neural networks (Valle-Perez et al., ), we include the geometric series penalizations $\lambda^{i}$ with $10<\lambda<$ 1 multiplying the reconstruction loss of length-$i$ initial subword of $m=S_{\phi}(x)$, and we gradually increase the value of $\lambda$ during training epochs. This process, a simple version of simulated annealing (Kirkpatrick et al., 1983), favors the model fitting the initial words of the messages at the beginning of training, and then gradually improving over the later words of the messages, as the coefficients $\lambda^{i}$ multiplying their contributions to the loss get higher.
Appendix D Pooling Examples
In this section we present two illustrative examples related to the statistics of pooling equilibria. We restrict to the simpler case of $\mathcal{G}=\mathcal{X}=\mathcal{D}$ and $\mathsf{GenX}=id$.
We say that an encoding is locally unambiguous (synonymous to being a pooling solution to the LSG) if it is decodable (i.e. injective) over the training set $\mathcal{D}_{train}$, and globally unambiguous if it is unambiguous over the whole of $\mathcal{D}$ (satisfies UE).
The two considered examples involve training sets $|\mathcal{D}_{train}|$ that differ by a factor of $2$, but they show that the proportion of numbers of global equilibria amongst local equilibria (i.e. the proportion of solutions to the UE condition over $\mathcal{D}_{train}$ that are also UE for $\mathcal{D}$) changes much more drastically, by a factor of $>10$ in this case.
Given a data distribution parameterized by three underlying factors with two values each, specifically let:
[…]
Example 1: easier compositional task. We will take a subset of the distribution as training data, and leave 2 configurations for evaluation, similar to an extrapolation problem (Schott et al., 2022). We will allow a channel capacity of 3 (i.e. take $C=3$) and a binary vocabulary:
[…]
Even though not all combinations are present in the training set, all generating factors are present. This situation becomes very likely when $|\mathcal{D}|$ becomes large, as proved in §A (see the simpler Prop. A.2 or equivalently Thm. A.3 with $k=1$).
We will define a encoding $S$ as a mapping from $\mathcal{X}$ (here $=\mathcal{D}$) to $[V]^{C}$ and a decoding $R$ as a mapping from $[V]^{C}$ to $\mathcal{D}$.
If we do not take into account decodability, encoding amounts to assigning some string in $\mathcal{M}=[V]^{C}$ to each state in $\mathcal{X}=\mathcal{D}$, there are $|\mathcal{D}|^{|V|^{C}}=8^{8}=16777216$ posible encodings.
For an encoding to be decodable without loss, $S$-encoding must form unique pairs from $\mathcal{D}$ to $[V]^{C}$ (condition UE in our setting). For local unambiguity it suffices that all training examples form unique pairs. Then, the amount of codings that are locally unambiguous is:
[…]
The amount of unique pairs from $\mathcal{D}_{train}$ to $[V]^{C}$ is the amount of injective mappings from $\mathcal{D}_{train}$ to $[V]^{C}$, i.e, $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}_{train}|)!}=20160$. The amount of encodings from $\mathcal{D}\setminus\mathcal{D}_{train}$ to $[V]^{C}$ is $(|\mathcal{D}|-|\mathcal{D}_{train}|)^{|V|^{C}}=2^{8}=256$. Therefore the amount of locally optimal encodings is $20160*256=5160960$. The number of globally optimal encodings is $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}|)!}=40320$. This implies that for this problem, the ratio of globally unambiguous to locally unambiguous solutions is $7,8125\times 10^{-3}$.
Example 2: hard compositional task. For the same $\mathcal{D}$ as above, we take
$\mathcal{D}_{train}:=\{\{\square,\rightarrow,\uparrow\},\{\circ,\leftarrow,\downarrow\}\},$
maintaining $C=3$ and $V=\{0,1\}$ as in the previous example.
This is, in a sense, ”the hardest possible” amongst solvable compositional reconstruction tasks, in the sense that this choice of $\mathcal{D}_{train}$ has minimum cardinality amongst training set selections such that all factor values are exemplified in some of their elements.
Compared to the previous example, the set of possible encodings and the set of globally unambiguous encodings remain the same, however, the amount of locally unambiguous mappings increases drastically compared to the previous example, and now is $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}_{train}|)!}*|\mathcal{D}\setminus\mathcal{D}_{train}|^{|V|^{C}}=94058496$, and thus the ratio of globally to locally unambiguous encodings decreases drastically to $4,28669\times 10^{-4}$
This illustrates the phenomenon that the ratio of global optima vs pooling solutions is very sensitive to the richness of $\mathcal{D}_{train}$.
Appendix E Hyperparameter search
Beta Lambda $\ell_{10^{-3}}$ Top Sim $\uparrow$ Final Symbol Loss $\downarrow$ Test Top Sim $\uparrow$ Test Final Symbol Loss $\downarrow$
0.5 0.75 4.7076 0.1802 0.0277 0.2333 0.0425
1 0.75 4.8052 0.1547 0.0512 0.2152 0.0520
10 0.75 4.9126 0.1597 0.0440 0.2156 0.0534
  0.5 0.95 5.2227 0.1894 0.0306 0.2681 0.0391
  1 0.95 5.3032 0.1761 0.0359 0.2335 0.0457
10 0.95 5.5674 0.1829 0.0436 0.2382 0.0456
0.5 1 5.3263 0.1766 0.0348 0.2526 0.0385
1 1 5.4198 0.1706 0.0362 0.2357 0.0492
10 1 5.5297 0.1661 0.0312 0.2284 0.0534
Table 2: Comparison of Hyperparameter Combinations (averaged over 10 runs), with first three metrics calculated at epoch 200 on training set, and the last two calculated at test after 200-epoch training. Recall that $\ell_{\epsilon}$ is the minimum message length, averaged over the dataset, here used with thresholding value $\epsilon=10^{-3}$. While low values of $\ell_{\epsilon}$ are better towards good compression (EE), this is only true assuming good accuracy, and otherwise small $\ell_{\epsilon}$ values may be unfavorable. TopSim abbreviates topological similarity. We highlight the combination used in the graphs in the body of the paper, $\beta=0.5$, $\lambda=0.95$, which, although not optimal for all metrics, is optimal for TopSim, and a good compromise on the other metrics.
Appendix F Generation examples
F.1 Reconstruction
The following reconstruction examples were generated with the ILFI+PDM model trained for 200 epochs, with $\lambda=0.95$ and $\beta=0.5$. From left to right and top to bottom, the first two images of each example are the original image to be reconstructed, and the reconstruction generated by the VAE backbone. The following images are the reconstructions produced by evaluated the predicted vector $z_{i}=R_{\omega}(S_{\phi}(x)_{[i]})$ with the VAE backbone.
We observe semantic differences between each prediction (flor hue, shape, object color, angle, etc.), as opposed to increasing in image quality or clarity. This fact leads us to believe that the message space effectively captures the generating factors $\mathcal{G}$ and is likely to be compositional in some measure.
Figure 4: Reconstructed images.
F.2 Interpolation
A possible way to evaluate the compositionality within the learned message space is to interpolate between representations of different objects in this space. Specifically, given two states $x_{1}$ and $x_{2}$, we visualize iteratively converting the message from one into the other.
In the following images, the first row consists of the original states $x_{1}$ (left) and $x_{2}$ (right). From the second row onwards, If $m^{1}=S_{\phi}(x_{1})$ and $m^{2}=S_{\phi}(x_{2})$, the $i_{th}$ row is the reconstruction from the message formed by concatenating $m^{1}_{[1:C-i]}$ and $m^{2}_{[C-i:C]}$, starting from 0 (i.e the second row from the top is simply $R_{\omega}(m^{1})$). We find that, as opposed to smoothly interpolating in generative model latent spaces, this exercise produces interpretable discrete steps to semantically interpolate the states. We invite the reader to observe the diagonal (starting from the rightmost cell on the third row), as this is where the message changes, and make note that the position of the diagonal coincides with some semantic change in the reconstruction.
Figure 5: Semantic interpolation 1.
Figure 6: Semantic interpolation 2.
Figure 7: Semantic interpolation 3.
Figure 8: Semantic interpolation 4.
Figure 9: Semantic interpolation 5.
Figure 10: Semantic interpolation 6.