### v4

# A Compressive-Expressive Communication Framework for Compositional Representations

Abstract
Compositionality in knowledge and language—the ability to represent complex concepts as a combination of simpler ones—is a hallmark of human cognition and communication. Despite recent advances, deep neural networks still struggle to acquire this property reliably. Neural models for emergent communication look to endow artificial agents with compositional language by simulating the pressures that form human language. In this work, we introduce CELEBI [^1] Our oficial implementation can be found at https://github.com/SugarFreeManatee/CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning), a novel self-supervised framework for inducing compositional representations through a reconstruction-based communication game between a sender and a receiver. Building on theories of language emergence and the iterated learning framework, we integrate three mechanisms that jointly promote compressibility, expressivity, and efficiency in the emergent language. First, Progressive Decoding incentivizes intermediate reasoning by requiring the receiver to produce partial reconstructions after each symbol. Second, Final-State Imitation trains successive generations of agents to imitate reconstructions rather than messages, enforcing a tighter communication bottleneck. Third, Pairwise Distance Maximization regularizes message diversity by encouraging high distances between messages, with formal links to entropy maximization. Our method significantly improves both the efficiency and compositionality of the learned messages on the Shapes3D and MPI3D datasets, surpassing prior discrete communication frameworks in both reconstruction accuracy and topographic similarity. This work provides new theoretical and empirical evidence for the emergence of structured, generalizable communication protocols from simplicity-based inductive biases.
1 Introduction
In natural languages, compositionality enables humans to communicate an infinite number of ideas using a finite set of elements [chomsky2014aspects-of-syntax]. This principle enables speakers to flexibly combine known words and structures to convey novel meanings, supporting flexible and generalizable communication across diverse and previously unseen contexts.
Works on the emergence of compositionality in language [iteratedlearningkirby, Kirby2002, KIRBY201587] have argued that opposing pressures are necessary for the natural selection of compositional languages between generations of speakers [elmoznino2024complexitybasedtheorycompositionality, beppu2009iterated]. On the one hand, successful communication requires high expressivity in order to usefully describe the world, allowing speakers to produce distinct messages for a wide range of meanings. On the other hand, models of emergent communication such as the iterated learning framework [iteratedlearningkirby] state that natural speakers tend to minimize the complexity of languages through cultural transmission, implying that simpler and more compressive languages are more easily passed on to new speakers.
These opposite pressures of expressivity and compressibility thus generate a trade-off, highlighted in several works [Kirby2002, elmoznino2024complexitybasedtheorycompositionality, beppu2009iterated] which is argued to be optimized by compositional languages, whereby languages evolve to maximize communicative efficiency – remaining expressive enough to convey diverse meanings while being simple and structured enough to be easily learned and transmitted by successive generations of speakers. When either pressure is dominant, undesirable languages tend to emerge: (a) excessive language compression leads to degenerate languages where multiple meanings are mapped to the same messages making them ambiguous and thus hard to use, and (b) prioritizing expressivity alone produces in holistic languages i.e. languages where messages are mapped arbitrarily to meanings without respecting their structures, hindering their transmission across generations [ren2020Compositional].
Enabling machines to generalize compositionally is thought to be crucial for them to quickly adapt to novel situations beyond their training experience [lake2017building]. Inducing compositional behavior in neural networks remains a major challenge [lake2017original-scan, keysers2019cfq, ruis2020gscan, kim2020cogs]. A growing body of work suggests that a model’s ability to generalize compositionally is highly sensitive to its training conditions [lake2019meta-seq2seq, conklin2021meta, ruis2022improving, lake2023human, delrio2025data], including factors such as data distribution, learning objectives, and task design. One promising direction is the study of emergent communication, where discrete languages evolve for coordination between independent neural agents.
In this work, we build on the Lewis reconstruction game framework [rita2022emergent] to study emergent communication in a reconstruction task. A sender encodes an image into a discrete message and a receiver reconstructs the original. Drawing on theories of language evolution [iteratedlearningkirby, Kirby2002], we develop a novel framework for inducing compositional communication grounded in simplicity bias.
We introduce CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning). Within this framework, we introduce three mechanisms in the learned communication protocol:
Progressive Decoding, in which the receiver makes reconstruction predictions after each incoming symbol, rather than waiting until the full message is received. This biases the system towards using intermediate reasoning steps, thereby imposing a pressure towards lower-complexity and less holistic encodings. While this mechanism does not directly improve reconstruction accuracy, it yields more efficient and structured communication.
Final-State Imitation modifies the standard imitation phase in the iterated learning (IL) framework of cultural evolution [Kirby2002, kirby2014iterated, KIRBY201587]. Instead of imitating the entire message, the student is trained to reproduce only the final output of the receiver, effectively reducing the information transmitted across generations. This tighter generational bottleneck increases pressure for compressibility in the emergent communication protocol [smithiterated], thereby promoting the emergence of compositional structure as a necessary condition for successful transmission. Empirically, this leads to increased compositionality with minimal degradation in reconstruction accuracy.
Pairwise Distance Maximization as a regularization term, i.e., as a further pressure towards increased diversity in the emergent language, that encourages exploration during the imitation phase. This term pushes the student sender to maximize an approximate Hamming distance between messages in a given batch, finding the most "diverse" protocol. We prove that this regularization gives a lower bound for entropy maximization [entropymax], and an upper bound on the contrastive learning loss NT-Xtent [chen2020simpleframeworkcontrastivelearning].
The opposing pressures induced by these three mechanisms impose a tight regularization on the complexity of the emergent language. These ingredients make explicit the pressures proposed in the cognitive science and linguistic literature to shape language evolution, thus aligning with the IL paradigm and leading to measurably higher compositionality.
2 Background
2.1 Problem setup: recovering compositional representations
[…]
The factors $\mathbf{G}$ are assumed to have compositional structure of the form $\mathbf{G}=[G_{1},\dots,G_{n}]$, in which $G_{i}$ represent independent characteristics having finitely many possible values (as in the dataset of Sec. 4.1 for example). In this case, an efficient way to reproduce elements of $\mathcal{D}$ is to impose that $\widehat{x}$ is similarly generated as $\widehat{x}=\mathsf{GenX^{\prime}}(\mathbf{G^{\prime}})$, with generating factors $\mathbf{G^{\prime}}$ giving a learned encoding of $\mathbf{G}$.
As $\mathcal{G}$ has a large number of classes (one per combination of factor values) and we have a single example per class, successful reconstruction relies on the compositional structure of $\mathbf{G}$. Thus the goal is to build a framework for finding the optimized compositional encoding $\mathbf{G^{\prime}}$ from observation of a small training set $\mathcal{D}_{train}\subset\mathcal{D}$. The compositional nature of $\mathbf{G}$ is what makes reconstruction possible from a small random $\mathcal{D}_{train}$, i.e. from data corresponding to a small subset of classes. This would be of course impossible for non-compositional data (see §D for a formal treatment).
2.2 Approach: Lewis reconstruction game
We frame the above reconstruction problem in the form of a cooperative game between two agents: a sender $S$ and a receiver $R$. The goal of $S$ will be to build the reconstruction factors $\mathbf{G}^{\prime}$ and the receiver $R$ will map the factors to an image $\widehat{x}$. The two agents aim to create a protocol that reconstructs $x$ to good accuracy. This setup fits in the general class of Lewis Reconstruction Games (LRG) [rita2022emergent], a subclass of so-called Lewis Signaling Games (LSGs) [lewis1969convention].
Within the framework of signaling games, the factors $\mathbf{G}^{\prime}$ are interpreted as a language which $S$ uses to communicate $x$ to $R$ as accurately as possible. The allowed message space is finite, i.e. $\mathbf{G^{\prime}}\in\mathcal{M}:=V^{C}$, and we take $C$ fixed and large enough that $|V|^{C}\gg|\mathcal{G}|$, so that the size of the message space is not a limiting factor on the emergent communication. The following diagram summarizes our notations:
$\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{D}\subseteq\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}=V^{C}\stackrel{{\scriptstyle R}}{{\longrightarrow}}\mathcal{X}$
Our main focus is on defining a IL process which favors the emergence of a compositional language between $S$ and $R$ that accurately recovers $x\in\mathcal{D}$.
2.3 Base model architecture
We adopt a standard emergent communication setup as described by elvae, see implementation details in § B.1. Sender and Receiver are implemented using the EGG framework [kharitonov:etal:2021]. Input images are encoded into latent representations using a pretrained VAE visual backbone. The sender encodes this latent input into a discrete message using an LSTM and then transmits it through a fixed-length communication channel. Message generation is made differentiable via a Gumbel-Softmax bottleneck [jang2017categoricalreparameterizationgumbelsoftmax]. The receiver takes the full message as inputs in an LSTM, and produces a latent vector, which is subsequently decoded into a reconstructed image by the VAE decoder. Reconstruction loss is computed between the original and generated images using Mean Squared Error (MSE). Both VAE encoder and decoder remain frozen during training.
3 Methods
This section describes our proposed improvements to the EL framework [elvae] (cf. Figure 1(a)). Intuitively, our additions attempt to increase the language drift between iterations of speakers, in order to thoroughly explore the landscape of possible languages, while maintaining efficient and useful communication. For the latter, we introduce Progressive Decoding (PD), to serve as an anchor for the emergent language during interaction, and for the former we introduce Final-State Imitation (FiSI) and Parwise Distance Maximization (PDM), which aim to incentivize exploration during the imitation phase. We expect the combination of these factors to show the strongest impact towards compositionality in the learned language, more so than the factors taken separately.
3.1 Progressive Decoding (PD)
Our first proposed enhancement to the baseline communication protocol aims to improve the efficiency of the emergent language determined by the Sender and Receiver, by condensing information in fewer message tokens. This type of efficiency has been argued to be necessary for the emergence of natural language properties [Kanwal2017-hs, Luna2020InternalAE].
Standard reconstruction games typically optimize a reconstruction loss computed only at the end of the communication phase, after the receiver has observed the full message [rita2022emergent], and as shown by rita-etal-2020-lazimpa, without an incentive to efficiently use the information received in the message, the receiver only updates its internal state after the full message has been received. To improve communication efficiency, we propose the following Progressive Decoding (PD) objective: the receiver generates a prediction after each received symbol, and our interaction phase loss explicitly weights the reconstruction error of each sub-message, as follows.
$\mathcal{L}^{\text{int}}_{\omega,\phi}:=\mathbb{E}_{x}\left[\frac{1}{C}\sum_{i=1}^{C}\lambda^{i}\,\text{MSE}\left(x,R_{\omega}(S_{\phi}(x)_{[i]})\right)\right],$ (1)
where $S_{\phi}(x)_{[i]}$ is the prefix of length $i$ of the sender’s message, and $\lambda\geq 1$ is an expressivity hyperparameter that increases the weight of later reconstructions. This loss has the following effects:
Efficiency pressure: Loss (1) rewards accurate reconstructions made as early as possible, favoring messages with shorter useful length.
Interpretable sub-messages: The model is encouraged to structure messages such that an interpretation by $R_{\omega}$ is available for sub-messages, making each symbol accountable for partial reconstruction. This echoes findings in prior theoretical and empirical work on chain-of-thought and compositionality [velickovic2020Neural, jiang2021inducing, chakravarthy2022systematicity, nye2022show, wei2022chain, prystawski2024think], which show that forcing intermediate semantic consistency or intermediate reasoning steps encourages models to show increased compositional behavior.
Tuning the efficiency pressure: Small $\lambda$ yields fast but potentially coarse reconstructions, while larger values favor detailed outputs at the cost of longer, possibly redundant messages, and in the limit $\lambda\to\infty$ (1) becomes equivalent to using just the last term in brackets, i.e., the full-length message reconstruction. In our experiments (see Section 5.2), we explore this trade-off empirically and show that $\lambda\equiv 1.5$ achieves the best performance across generalization, communication efficiency metrics, and compositionality.
3.2 Imitation Phase
Iterated learning is known to add compressibility pressure to emergent communication schemes [iteratedlearningkirby], resulting in languages that are "easy to learn", thus favoring simple communication schemes that can be compressed: this is known as a communication (or generational) bottleneck. It is argued [iteratedlearningkirby] that languages that remain relatively stable even when a learner only observes a small subset of the language of the previous generation need to be compositional.
Our case (and in fact, most deep learning applications of IL) differ from the original IL formulation [iteratedlearningkirby], in that the transmission bottleneck does not constrict the amount of examples presented to the student, but rather constricts the information obtainable from these examples in order to perfectly reconstruct the language. This paradigm shift for transmission bottlenecks is at the heart of discrete bottlenecks [ren2020Compositional, rita2022emergent, rita-etal-2020-lazimpa] and noisy bottlenecks [ren2024improving].
We next pass to describe our proposed imitation scheme and its motivations compared to previous methods.
3.2.1 Final-State Imitation (FiSI)
As usual in IL, we apply a teacher-student regime to initialize the next iteration sender $S_{\phi^{t+1}}$.
Our main proposed novelties for the imitation phase are as follows.
State space imitation: As opposed to other IL implementations [ren2020Compositional, ren2020Compositional, li2019easeofteachinglanguagestructureemergent], our models are not directed to match the student protocols from different iterations in the message space, but rather in the state space, using the frozen receiver network from the interaction phase. This change allows a wider range of sender strategies, as the new loss is more permissive.
Final state reconstruction: During imitation phase we use the following reconstruction loss, where $R^{\star}_{t}=R_{\omega^{t}},S^{\star}_{t}=S_{\phi^{t}}$ are frozen receiver and sender from the previous iteration and $S_{t+1}=S_{\phi^{t+1}}$ is the newly trained sender:
$\mathcal{L}^{rec}_{\phi^{t+1}}:=\mathbb{E}_{x}\left[d_{\mathcal{X}}(R^{\star}_{t}(S^{\star}_{t}(x)),R^{\star}_{t}(S_{\phi^{t+1}}(x)))\right].$ (2)
Our loss $\mathcal{L}^{rec}_{\phi^{t+1}}$ only depends on the state associated to the full message, without testing reconstruction on sub-messages. This change allows the intermediate message tokens of the student sender $S_{\phi^{t+1}}(x)_{i},i<C$ to drift further away from the ones of the teacher mapping $S^{\star}_{t}(x)_{i}$ and thus more easily find different strategies with the same end result.
State space imitation is justified as follows. Note that as message space is discrete, there exists a value $\delta_{0}$ such that any two messages at distance $\delta_{0}$ in fact coincide. Then $d_{\mathcal{M}}(S^{\star}_{t}(x),S_{t+1}(x))<\delta_{0}$ always implies $d_{\mathcal{X}}(R^{\star}_{t}(S^{\star}_{t}(x)),R^{\star}_{t}(S_{t+1}(x)))=0$, while the vice-versa is not always true: if $R^{\star}_{t}$ is not injective, any $S_{t+1}(x)\in(R^{\star}_{t})^{-1}(S^{\star}_{t}(x))$ will preserve zero distance of $R^{\star}_{t}$-images. If $R^{\star}_{t}$ has mild Lipschitz regularity assumption, this principle extends to the regime of small loss, showing that our new loss is less restrictive on the choices of $S_{t+1}$ than the traditional message space loss. Thus, state space imitation enables wider sender strategy exploration without penalization losses.
(a) Overview of our proposed architecture
(b) Qualitative example of image decoding
Figure 1: (a) Overview of our proposed architecture. Interaction Phase: The sender $S_{\phi^{t}}$ and receiver $R_{\omega^{t}}$ are jointly trained to minimize the reconstruction error between the input state $x$ and the predicted states $\{x^{\prime}\}$, by encoding $x$ into a message $m^{t}_{x}$. Imitation Phase: A new sender $S_{\phi^{t+1}}$ is trained to imitate the final predicted output of $R_{\omega^{t}}(m^{t}_{x})$, while also maximizing pairwise message diversity to encourage exploration. (b) Qualitative example of image decoding: the receiver reconstructs the input image from the sender’s message for each sub-message, giving a series of reconstructions. The useful length of a message corresponds to the first reconstruction that has error below a threshold $\epsilon>0$: in this example the last two message tokens are not useful in that they do not add to the reconstruction accuracy.
The choice of using just final states is justified because final state reconstruction allows greater freedom during imitation for the student sender. By only constraining the reconstruction of the full message, we allow much more freedom for the encoding strategy of $S(x)$ at earlier sub-messages at each iteration (see Proposition E.1 for a formal statement and proof).
3.2.2 Pairwise Distance Maximization (PDM)
To further improve protocol diversity and promote exploration of the language space during the imitation phase, we introduce a regularization objective that promotes dissimilarity between messages within each batch. In addition to enabling a larger search space for optimal student strategies. Enhancing the diversity of student strategies also fits the general IL argument according to which compositional languages are the most resistant to noisy language transmission [iteratedlearningkirby, KIRBY201587, kirby2014iterated].
Specifically, we approximate the Hamming distance by computing the position-wise cosine dissimilarity between the probability distributions over symbols in each message. [^2] For one-hot encoded messages, this formulation is equivalent to maximizing the Hamming distance. This encourages the sender to generate messages that are maximally distinct while remaining semantically aligned with the teacher protocol. The strength of this regularization is controlled by a hyperparameter $\beta$.
The resulting imitation loss is defined as:
$\mathcal{L}^{im}_{\phi^{t+1}}:=\mathcal{L}^{rec}_{\phi^{t+1}}+\beta\frac{1}{N_{batch}^{2}}\sum_{i,j}d(S_{\phi^{t+1}}(x_{i}),S_{\phi^{t+1}}(x_{j})),$ (3)
where $d(\cdot,\cdot)$ is the mean cosine similarity between corresponding symbol distributions.
In § F.4, we show that this regularization term provides a lower bound on entropy maximization objectives [entropymax], and serves as an upper bound for contrastive losses such as NT-Xent [chen2020simpleframeworkcontrastivelearning], thus connecting our formulation to well-established principles in representation learning.
4 Experimental setup
4.1 Datasets
Shapes3D
We tested our framework using the Shapes3D [dsprites17] dataset, which consists of colored images of 3D geometric shapes, with 6 underlying generating factors $G$: floor hue, wall hue, object hue, shape, scale, and orientation. The total amount of attribute-value combinations is $480,000$.
MPI3D
We also evaluated using the compositional split of the MPI3D dataset [gondal2019transfer]. This dataset consists of colored images of a robot arm interacting with objects, rendered in controlled 3D scenes with 7 underlying generative factors $G$: object color, object shape, object size, camera height, background color, horizontal arm position, and vertical arm position. The dataset comprises $1,036,800$ unique combinations of these factor values.
For both datasets we use the compositional split from schott2022visual, which ensures all attribute values appear in training, yet some combinations are reserved for the testing set (See §D for a proof that this situation would anyways hold for random train/test splittings with high probability).
4.2 Evaluation Metrics
Following chaabouni2022emergent, we assess compositionality using Topographic Similarity (TopSim) [Brighton2006UnderstandingLE], a widely used proxy for compositionality. TopSim measures the correlation between pairwise Hamming distances in message space and the generating factor space, capturing the extent to which semantically similar inputs yield similar messages under a structured encoding.
To evaluate communication efficiency, we define the useful length $\hat{\ell}_{\epsilon}(x)$ , which estimates the minimal prefix of a message necessary to achieve near-maximal reconstruction quality. Formally:
$\hat{\ell}_{\epsilon}(x):=\min\left\{i\in\{1,\dots,C\}:\text{MSE}\left(x,R_{\omega}(S_{\phi}(x)_{[i]})\right)\leq\epsilon\right\},$
where $C$ is the maximum message length and $\ell_{\epsilon}(x)=C$ if the threshold is unmet.
We choose $\epsilon$ by viewing the loss distributions on both tested datasets and estimating a common plateau point for each. The full position-wise loss values can be found in the supplementary material.
Finally, to approximate the expressivity of the communication protocol, we evaluate reconstruction quality using the MSE between the generated image and the original input.
We discuss the use of additional language metrics in §I.
5 Results
For all experiments, we report the mean and standard error of 10 random seeds. To assess the robustness of our findings, we also perform permutation tests to evaluate the statistical significance of the observed differences (see §J.1). Additionally, we include qualitative reconstruction experiments in §H, where we visually compare the outputs of all evaluated methods across progressive decoding steps. Further implementation details are provided in §B.
5.1 Progressive Decoding reduces the message’s useful length
Table 1: Performance comparison across all experiments on the Shapes3D and MPI3D datasets. We report Topographic Similarity (TopSim $\uparrow$), Useful Length ($\hat{\ell}_{\epsilon}$ $\downarrow$), and Last Symbol MSE ($\downarrow$) for each ablation. The results are grouped by the components of our proposed framework: Interaction, Imitation, and Regularization. The synergetic mixture of our proposed methods (PD+FiSI+PDM) consistently improves compositionality and message efficiency over the baseline and prior variants.
TopSim $\uparrow$ $\hat{\ell}_{2\times 10^{-1}}$ $\downarrow$ Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.244\pm 0.002$ $10.0\pm 0.0$ $0.212\pm 0.003$
Progressive Decoding(PD, ours) $0.270\pm 0.001$ $7.5\pm 0.215$ $0.238\pm 0.007$
Imitation Message imitation (baseline) $0.257\pm 0.005$ $9.9\pm 0.3$ $0.200\pm 0.007$
Full state imitation $0.256\pm 0.003$ $9.9\pm 0.3$ $0.179\pm 0.003$
Final-State Imitation (FiSI, ours) $0.283\pm 0.003$ $10.0\pm 0.0$ $\mathbf{0.176\pm 0.002}$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.256\pm 0.002$ $8.555\pm 0.167$ $0.179\pm 0.003$
PD+FiSI+PDM (ours) $\lambda=1.5$ $\mathbf{0.292\pm 0.002}$ $\mathbf{7.0\pm 0.155}$ $0.194\pm 0.005$
(a) Shapes3D
TopSim $\uparrow$ $\hat{\ell}_{1.95\times 10^{-2}}$ $\downarrow$ Last symbol MSE $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.133\pm 0.001$ $9.3\pm 0.002$ $\mathbf{0.015\pm 0.0}$
Progressive Decoding(PD, ours) $0.137\pm 0.001$ $\mathbf{6.6\pm 0.341}$ $\mathbf{0.015\pm 0.0}$
Imitation Message imitation (baseline) $0.135\pm 0.001$ $9.733\pm 0.029$ $0.016\pm 0.0$
Full state imitation $0.137\pm 0.001$ $9.923\pm 0.020$ $0.018\pm 0.0$
Final-State Imitation (FiSI, ours) $\mathbf{0.156\pm 0.001}$ $9.7\pm 0.046$ $0.02\pm 0.0$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.147\pm 0.002$ $8.9\pm 0.221$ $0.02\pm 0.0$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.153\pm 0.001$ $9.0\pm 0.167$ $0.02\pm 0.0$
(b) MPI3D
We begin by evaluating the effect of incorporating PD during the interactive phase of training. As shown in the first two rows of Table 1(b), this modification leads to a 25-29% reduction in the useful message length, indicating more efficient communication. These gains are achieved without compromising expressivity on MPI3D and with only a slight compromise on Shapes3D, as indicated by stable reconstruction quality in the former and a minor degradation in the latter.
To further promote efficient and structured communication, we introduce a geometric penalty term $\lambda$ that penalizes reconstruction error more heavily when longer messages are used. As shown in Figure 2, increasing $\lambda$ up to values near 1.5 consistently improves reconstruction quality and useful length, indicating higher expressivity and efficiency. However, when $\lambda$ becomes too large, the reconstruction error incurred at the end of the message dominates the loss function, effectively nullifying PD. We found that setting $\lambda=1.5$ yields favorable results in efficiency and expressivity, while maintaining high compressibility.
(a) Useful Length
(b) Reconstruction MSE
(c) TopSim
Figure 2: Effect of geometric penalty $\lambda$ on emergent communication. We introduce a geometric weighting term $\lambda$ that penalizes reconstruction error more strongly for longer messages. Increasing $\lambda$ initially improves useful length (a). However, once a certain threshold is met, useful length and reconstruction error (b) increase. Values close to 1.5 of $\lambda$ strike a favorable balance—achieving efficient and expressive communication.
5.2 Final State Imitation increases TopSim
Second, we investigate the impact of final state imitation on the structure of the emergent language.
In these experiments, we compare our approach to a baseline model without imitation and to a standard IL variant that imitates the teacher’s message directly—a strategy commonly adopted in prior work for optimizing this phase [ren2020Compositional, ren2024improving, li2019easeofteachinglanguagestructureemergent].
As shown in the imitation section in Table 1(b), in the MPI3D dataset, we find that our final state imitation method consistently enhances the compositionality of the emergent messages. It outperforms the non-IL baseline and surpasses message-level imitation significantly (§ J.1).
More broadly, our results support prior findings that IL promotes greater structure in emergent communication protocols [li2019easeofteachinglanguagestructureemergent, ren2020Compositional]. They also underscore the value of shifting the learning objective from message imitation to image reconstruction, which enforces a tighter generational bottleneck by requiring the student to recover the intended output without access to the original message.
Moreover, this setup increases flexibility in protocol exploration by constraining the student only to the reconstruction target, not the exact message form. Permitting generational variation in message structure while preserving semantic fidelity, promoting diversity without sacrificing expressivity.
Together, these factors exert pressure toward discovering more systematic and compositional protocols—structures that are inherently more transmissible and robust across generations.
5.3 Pairwise Distance Maximization increases compositionality and efficiency
We evaluate a model that combines all previously identified best-performing components—specifically, $\lambda=1.5$, final-state imitation, PD and further introduce an entropy-based regularization term using PDM. We compare this with the KoLeo entropy estimator proposed by sablayrolles2018spreading.
When examining the effect of incorporating PDM, Table 1(b) shows that PDM outperforms both FiSi and PD in fostering structure in the message. On the Shapes3D dataset, PDM leads to a marked improvement in TopSim and a lower useful message length, indicating more efficient and structured communication, albeit with a slight increase in reconstruction error. For MPI3D, this configuration doesn’t achieve the highest TopSim, albeit close to the best, and it increases useful length and leads to a marginally higher reconstruction loss compared to using only PD. This highlights a trade-off between expressivity and compressibility, in line with prior theoretical accounts on communication efficiency and compositionality [ambiguityandefficiency, iteratedlearningkirby, Luna2020InternalAE].
When comparing PDM to the KoLeo estimator, Regularization section in Table 1(b), we find that on Shapes3D, PDM achieves superior TopSim and shorter useful length, with slight reconstruction gains. On MPI3D, KoLeo leads to a lower useful length but exhibits reduced compositionality, with both regularizers showing similar reconstruction performance.
5.4 Comparison to disentangled representation learning
Following the comparison made by elvae, we test our generated messages against standard self-supervised disentangled representation learning frameworks, namely $\beta$-VAE and $\beta$-TCVAE, as well as the baseline VAE+EL [elvae]. For the continuous models, we extract the first half of the latent vector (corresponding to the predicted means $\mu$) and evaluate it using the DCI disentanglement score [DCI2018]. In addition, we train a two-layer MLP to predict ground-truth generative factors, reporting RMSE for continuous factors and classification accuracy for categorical ones.
For the discrete models, we use the predicted messages for each image. We calculate the DCI disentanglement as-is, and use an embedding matrix to transform the discrete messages into the same dimension as the continuous vectors for the MLP. Training is conducted on three subsets of size 1000 of the training set and averaged. Evaluations are conducted on the test sets of the compositional splits for the Shapes3D and MPI3D datasets.
Across both datasets, our method (VAE+CELEBI) demonstrates consistent improvements in disentanglement over baseline methods. On Shapes3D, we observe a substantial increase in DCI scores compared to continuous models and the discrete baseline (see statistical significance tests in § J.2). However, both discrete models fall behind the continuous baselines in downstream accuracy and RMSE, suggesting a trade-off between representation interpretability and usefulness in downstream tasks, which is consistent with previous work indicating that discrete representations may be less accessible to simple classifiers (as shown in [xu2022compositional]), and that inductive biases for compositionality do not imply disentanglement [montero2021the].
On MPI3D, our method slightly under-performs when compared to the continuous baselines in categorical accuracy, however greatly outperforming all baselines in continuous regression. We speculate that this discrepancy arises from the lower correlation between pixel-level input statistics and continuous generative factors in MPI3D as opposed to Shapes3D, which may favor communication-based models over purely reconstructive.
Table 2: Disentanglement metrics for the Shapes3D and MPI3D datasets. We report the DCI score to evaluate disentanglement quality, and the RMSE and classification accuracy of a linear probe trained on the latent representations (both message and vector embeddings) to assess their suitability for downstream tasks.
Dataset Model Disentanglement $\uparrow$ RMSE $\downarrow$ Acc $\uparrow$
Shapes3D $\beta$-VAE $0.045\pm 0.004$ $2.460\pm 0.144$ $0.702\pm 0.025$
$\beta$-TCVAE $0.043\pm 0.005$ $\mathbf{2.378\pm 0.123}$ $\mathbf{0.721\pm 0.028}$
VAE + EL $0.108\pm 0.000$ $3.168\pm 0.027$ $0.343\pm 0.019$
VAE + CELEBI $\mathbf{0.112\pm 0.001}$ $2.932\pm 0.045$ $0.453\pm 0.024$
MPI3D $\beta$-VAE $0.031\pm 0.001$ $21.819\pm 0.111$ $0.581\pm 0.006$
$\beta$-TCVAE $0.031\pm 0.001$ $21.854\pm 0.188$ $\mathbf{0.582\pm 0.005}$
VAE + EL $0.114\pm 0.002$ $15.971\pm 0.301$ $0.529\pm 0.006$
VAE + CELEBI $\mathbf{0.137\pm 0.002}$ $\mathbf{14.82\pm 0.012}$ $0.542\pm 0.002$
6 Related Work
The use of language reference games (LRGs) in conjunction with iterated learning for the purpose of learning self-supervised representations remains underexplored. elvae compare an emergent language autoencoder—comprising a visual backbone and an LSTM-based sender and receiver—with disentanglement frameworks such as the $\beta$-VAE and $\beta$-TCVAE. They find that the representations induced by the emergent language model generalize better in downstream tasks, highlighting the potential of symbolic communication as a compositional bottleneck.
Many previous works have explored using emergent language to induce compositional behavior [kharitonov2020entropy, Dess2021InterpretableAC, auersperger2022defendingcompositionalityemergentlanguages, Denamganai2020OnS, chaabouni2020compositionality]. Several works utilize a referential LSG in conjunction with IL, using predicted messages as a generational bottleneck [ren2020Compositional, li2019easeofteachinglanguagestructureemergent]. Alternative bottlenecks for emergent language have been proposed, including simplicial embeddings [ren2024improving, Elberg_2024], code-books [zheng2024iterated] and noisy channels [ueda2024lewisssignalinggamebetavae]. Our message tokens can also be considered as roughly comparable to the slots in slot-attention architectures such as [eslami2016attend, greff2019multi, Lin2020SPACE, jiangscalor, lin2020improving, Engelcke2020GENESIS, locatello2020object, brady2025interaction, akan2025slotguided], however note that among other differences, the slot-attention typically additive decoding layer [burgess2019monet, lachapelle2023additive] does not compare directly to our proposed Progressive Decoding.
To replicate the Zipfian distributions observed in natural languages [Zipf49, Kanwal2017-hs], rita-etal-2020-lazimpa propose the LazImpa framework. The LazImpa Impatient listener loss is similar to our proposed Progressive Decoding, in that they both aim to induce incremental informativeness and efficiency by reconstructing the input at every timestep. However LazImpa is optimized for a referential game task (and not a reconstruction task as ours) and uses a linear length penalization in conjunction with a cross-entropy loss. Our proposal instead uses the actual reconstruction loss and the length penalization is exponential and based on a tunable parameter $\lambda$.
7 Limitations and Future Work
Our experiments are conducted exclusively on the synthetic datasets Shapes3D and MPI3D [3dshapes18, gondal2019transfer], which feature clean, disentangled generative factors and lack observational noise. In future work, we plan to extend our the same principle of making explicit pressures towards complexity reduction of underlying representations to natural image datasets where factor supervision may be unavailable or weakly defined. Furthermore, we aim to formalize the notion of language compression defined in IL in terms of Kolmogorov complexity, for representing datasets with more complex compositional structures.
Moreover, it is debatable whether MSE is the most appropriate choice for the self-supervised loss function, especially in the context of natural image data. While MSE is a simple choice, future work may explore reconstruction losses that better align with human visual perception, such as perceptual similarity metrics based on deep feature embeddings (e.g., LPIPS [lpips]), or adversarial and contrastive losses that promote more semantically structured representations.
Additionally, we froze the visual backbone across all EL model variants trained on the same dataset. While this ensures controlled comparisons within each dataset, it may introduce an initialization bias when comparing with continuous baselines, where the encoder is jointly optimized. Although our results remain statistically significant, future work could investigate whether co-training the vision encoder yields improved alignment between symbolic and perceptual representations, or leads to new trade-offs in compression and expressivity.
Finally, our experiments were restricted to compositional generalization metrics and reconstruction-based evaluation. Other facets of emergent communication, such as robustness and interpretability, are left for future work. We believe our framework offers a strong foundation for such extensions and a tractable setting to explore the interplay between compression, efficiency, and generalization.
8 Conclusion
We introduce the CELEBI framework with three novel mechanisms for enhancing compositional learning in IL frameworks, as well as mathematical justification for their design. First, with Progressive Decoding (PD) we reward informative communication at each step of the message, creating an inductive bias toward efficient and distributed representations. We show that PD promotes both message compressibility and compositional alignment.
Second, in our Final State Imitation (FiSI) in IL, the student is trained to reproduce the final prediction of the teacher, rather than the message itself. This shift enables greater exploration of the message space while preserving semantic consistency. We provided theoretical motivation and empirical evidence that this approach yields more expressive and compositional protocols, outperforming traditional message imitation in TopSim and reconstruction metrics.
Third, our regularizer based on Pairwise Distance Maximization (PDM), which provably approximates entropy maximization over messages, serves as a practical inductive bias for promoting diversity and structure in emergent languages, particularly during the imitation phase.
Together, these contributions enrich the IL framework, and give new insights into how training dynamics and inductive pressures shape the emergence of language-like representations. As confirmed by our empirical findings, the proposed methods lead to more compositional and generalizable communication schemes.
Acknowledgments and Disclosure of Funding
This work was supported by ANID Chile, Fondecyt Regular grant 1231724, as well research centers of excellence with code FB210017 (Basal CENIA), ICN2021_004 (Millenium iHealth), and ICN17_002 (Millenium IMFD).
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims and contributions are clearly stated in the abstract and introduction, and are backed by theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper contains a separate limitations section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical results are rigorously proven either in the main text or in the supplementary material, and justifications of the result’s hypotheses are given.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All implementation details, including hyperparameters, used libraries and experimental details, are included either in the main text or in the supplementary material. We plan to release the all code upon acceptance.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our full code and details to reproduce our results can be found in the oficial repository.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We use and disclose publicly available splits of known datasets, and specify all training hyperparameters either in the main text or the supplementary material.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We disclose the amount of seeds used in each experiments, and show standard error for the results in all figures and tables.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details of the computational resources used for this paper in the Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the code of ethics and have confirmed that our research complies with every applicable requirement.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not foresee any significant societal impact, as the paper explores principles for compositional learning and not its applications to society.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve models with high risk of misuse, as it explores simplified models for the sake of finding new principles for enhancing compositionality of representations based on an iterated communication game setup.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All assets are properly credited and used in accordance with their licenses: Shapes3D and MPI3D are under CC-BY 4.0, and the EGG framework is released under the MIT license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not use crowdsourcing experiments or experiments with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The paper’s results were not based on, or aided by, research using LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Background
A.1 Iterated Learning
The IL model of cultural evolution [iteratedlearningkirby] proposes to emulate the emergence of language through the interactions of adult and new learning agents. In the original formulation, a space of pairs of signal + meaning is randomly generated, and in each iteration, a learning agent is partially exposed to the language learned by the adult agent in the previous iteration. The learning agent becomes the adult agent for the next iteration, and transmits a subset of its learned "language". By iterating this process, the initial random language gains structure and observable properties, such as compositionality.
This framework has been widely extended to include an additional task [li2019easeofteachinglanguagestructureemergent, ren2024improving, zhou2022fortuitousforgettingconnectionistnetworks], instead of simply reshaping the language. Under this formulation, the training regime can be separated into two parts: Interaction (Interaction), where an agent is optimized to solve the additional task, and Imitation (Imitation), where a newly initialized agent "learns" from the previous agent through a shared learned language.
A.2 Lewis reconstruction game
The Lewis reconstruction game (LRG) [rita2022emergent] is a special case of the Lewis Signaling Game (LSG) [lewis1969convention] framework, in which two agents cooperate to reconstruct an observed object. Specifically, a sender agent parameterized by $\varphi$ observes an object $x$ from an object space $\mathcal{X}$ and produces a message $m\in\mathcal{M}$, where $\mathcal{M}=V^{\star}$ is the set of all possible strings formed by concatenating symbols drawn from a finite vocabulary $V$. A receiver agent parameterized by $\omega$ observes $m$ and produces a reconstruction $x^{\prime}$. Both models are optimized to minimize a reconstruction loss $\mathcal{L}_{R}(x,x^{\prime})$.
Importantly, in the original LSG, the receiver is tasked with predicting $x$ from a finite set of states, whereas in this formulation, the problem becomes a regression task, and $\mathcal{X}$ can be a continuous space.
Appendix B Implementation Details & Hyperparameters
B.1 Baseline Model
  • Sender and Receiver: Both consist of a single-layer LSTM with an embedding size of 64 and a hidden size of 256. Outputs are passed through a two-layer MLP with ReLU activations.
  • Communication Channel: The vocabulary size is set to $|V|=15$ and messages are composed of $C=10$ discrete symbols. These values are chosen such that the amount of posible messages ($5.76\times 10^{1}2$) greatly outnumbers the amount of states. The Gumbel-Softmax bottleneck is used without straight-through estimation.
[…]
  • VAE Training: The VAE is pretrained for 15 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$.
The interaction-imitation games were run for a maximum of 100 iterations, alternating one full epoch in each phase per iteration.
We used early stopping based on the validation MSE of the final reconstruction, starting from epoch $5$, with a minimum $\delta$ of $1\times 10^{-3}$ and patience of $5$. We used the Adam optimizer with default parameters and learning rates of $1\times 10^{-3}$ and $1\times 10^{-4}$ for Shapes3D [3dshapes18] and MPI3D [gondal2019transfer], respectively, as we found the training for the latter became unstable at higher learning rates.
We found that Interaction required smaller batch sizes to converge, whereas the Imitation phase PDM becomes a better entropy approximation at higher batch sizes, however scaling quadratically in computation. Therefore, we set the Interaction and Imitation batch sizes to 256 and 512, respectively.
To implement Imitation we defined a receiver $R_{\omega}$ and two senders $S_{\phi_{1}}$ and $S_{\phi_{2}}$ with their own optimizers. After participating in an Interaction iteration, $\phi^{1}$ and $\omega$ are frozen and used to train $S_{\phi_{2}}$ as described in §3.2. After Imitation, the weights $\phi_{2}$ are copied to $\phi_{1}$, and $\phi_{2}$ is reset. Importantly, the optimizers for $\phi_{1}$ and $\phi_{2}$ are never reset nor copied between them.
Appendix C Compute Resources
All experiments were conducted on our internal laboratory cluster, using NVIDIA A40 GPUs with 48 GB of VRAM. Each job was allocated a single GPU with 8 GB of VRAM usage on average, alongside access to a CPU with 20 cores, 118 GB of RAM, and local SSD storage for datasets and model checkpoints.
The experiments reported in this paper comprise 360 training runs, totaling approximately 180 GPU hours. Individual runs varied in length from 15 to 50 minutes, depending on early stopping criteria. These runs represent the finalized experiments whose results are presented in the main text.
In total, the research project required 2,310 experimental runs over the course of development, amounting to approximately 245 GPU days (5,880 GPU hours) of compute time. This includes preliminary experiments, ablations, and failed runs that were instrumental to model and protocol design but are not individually reported in the paper.
C.1 Useful length $\epsilon$
To define the useful length $\epsilon$ for both datasets, we observed reconstruction loss distribution over all positions in the message (see 3) for different $\lambda$ values, and estimated an average plateau point. For MPI3D we used a more conservative and precise threshold, as the reconstruction error varied more between methods.
Figure 3: Loss distribution over positions and $\lambda$ values for (a) Shapes3D and (b) MPI3D
Figure 4: Useful length values for different $\epsilon$ and $\lambda$ values for (a) Shapes3D and (b) MPI3D
Additionally, we analyzed the useful length graph for multiple $\epsilon$ values, and discarded those values where most models converged to either maximum or minimum possible length (see 4).
Appendix D Recovering compositional structure from partially observed data
For the sake of clarity of treatment, we assume here that $\mathbf{G}=[G_{1},\dots,G_{n}]$ in whichi factor $G_{i}$ has the same number $N$ of possible values, and as a consequence $\mathcal{G}=[1:N]^{n}$. Our results will extend directly to the general $\mathcal{G}$ case as shown in Rmk. D.4.
We work under the simplified deterministic generation hypotheses, in which our dataset is $\mathcal{D}=\mathsf{GenX}(\mathcal{G})\subset\mathcal{X}$ in which $\mathsf{GenX}$ is a deterministic injective function.
Our main result is the following:
Theorem D.1.
Assume that $\mathcal{G}=[1:N]^{n}$ and that $\mathsf{GenX}:\mathcal{G}\to\mathcal{X}$ is an injective function, and that $\mathcal{D}_{train}\subseteq\mathcal{D}$ has cardinality $|\mathcal{D}_{train}|=p|\mathcal{D}|$ for some $p\in(0,1)$ such that $pN^{n}$ is an integer.
We assume that the compositional structure of $\mathcal{G}$ is such that the following holds:
  1. (A)
    For some $k<n$, if we observe a set of data whose generating factor combinations feature all possible combinations of generating factor values $(G_{i_{1}},\dots,G_{i_{k}})$ for all choices of indices $1\leq i_{1}<\cdots<i_{k}\leq n$ then this is sufficient to reconstruct $\mathcal{G}$ to good accuracy.
Then as $|\mathcal{G}|\to\infty$ the probability that $\mathcal{D}_{train}$ is sufficient to reconstruct $\mathcal{G}$ tends to $1$.
Here (A) makes rigorous the common assumption that having a training set that features examples for diverse enough combinations of factors is sufficient for reconstructing all combinatorial generators $\mathbf{G}$.
Note that in this work we do not discuss
  1. 1.
    what features of the pre-trained encodings $\mathcal{G}\to\mathcal{D}$ and what requirements on the compositional structure of $\mathcal{G}$ would allow to actually guarantee this assumption,
  2. 2.
    alternatives for a rigorous definition of what is meant by "to good accuracy" in the statement of assumption (A).
Once the above points are settled and defined, we can replace (A) by an explicit requirement. Since however these points seem like an ambitious whole research line and are far beyond the scope of this work, we leave the definition of "to good accuracy" and the proof of reconstruction to future works.
Thus, introducing (A) allows us to prove fully rigorous statements such as Thm. D.3, in absence of a full theory of compositional reconstruction.
Nevertheless, we make the following remarks:
  • There is a natural trade-off between assumptions on how big $\mathcal{D}_{train}$ should be, vs. how complicated the compositional reconstruction should be allowed to be for reconstruction to be feasible: the richer set of such combinations we require $\mathcal{D}_{train}$ to contain (thus making more restrictive assumptions on $\mathcal{D}_{train}$), the lower is the requirement for our encoding framework and for the compositional structure for (A) to hold.
  • It seems natural to conjecture that if assumption (A) does not hold for large value of $k$, then there is no hope for the compositional reconstruction to be achievable. Thus lowering the requirement on "good enough accuracy" to make (A) hold seems the only avenue of future research.
Theorem D.1 is a consequence of below Theorem D.3 that proves that assumption (A) holds with probability tending to $1$ as $|\mathcal{G}|\to\infty$ and $|\mathcal{D}_{train}|/|\mathcal{G}|=p$ is bounded away from zero. What Theorem D.3 says is that for large sets of generators, it will become overwhelmingly unlikely that $\mathcal{D}_{train}$ does not contain enough data to reconstruct the generators.
D.1 Observing $\mathcal{D}_{train}$ allows to observe finite statistics of the factors with high probability
Our first result for this section is that when $\mathcal{D}_{train}\subseteq\mathcal{D}$ is chosen uniformly at random amongst sets of size $p|\mathcal{D}|$, for $|\mathcal{D}|=|\mathcal{G}|$ large enough, we find that with overwhelming probability all possible values of each of the factors $G_{1},\dots,G_{n}$ are achieved by some elements of $\mathcal{D}_{train}$. We state the result in Proposition D.2, whose proof in our view contains the main ideas useful also for the full result of Theorem D.3, which can be considered our main technical result for this section.
The first result is as follows:
Proposition D.2.
Assume that for $n,N\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N\}=[1:N]^{n}$ and that $\mathsf{GenX}:\mathcal{G}\to\mathcal{X}$ is an injective function with image $\mathcal{D}$ as above. Denote by $\mathcal{G}_{train}:=\mathsf{GenX}^{-1}(\mathcal{D}_{train})$ the generating factors related to the training set, and assume that $\mathcal{D}_{train}\subset\mathcal{D}$ is chosen uniformly at random amongst subsets of cardinality $p|\mathcal{D}|$, for some $p\in(0,1)$ such that $pN^{n}$ is an integer. Then
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}\leq nN\left(1-\frac{1}{N}\right)^{pN^{n}},$ (4)
in particular at fixed $p$, the above event has probability tending to zero as $N^{n}\to\infty$.
Thus, in an equivalent formulation of the last part of the above proposition, we have that in the limit of large $|\mathcal{G}|=N^{n}$, the probability that each generating factor $G_{i}$ takes all of its possible values on some generated element of $\mathcal{G}_{train}$ tends to $1$.
Proof.
The possible choices of $\mathcal{G}_{train}\subseteq\mathcal{G}$ of cardinality $p|\mathcal{G}|=pN^{n}$ are $\tfrac{(N^{n})!}{(pN^{n})!((1-p)N^{n})!}$. Then, calling $A_{ij}$ the set of elements $\mathbf{G}\in\mathcal{G}_{train}$ such that $G_{i}\neq j$, we find that $|A_{ij}|=N^{n-1}(N-1)=N^{n}\tfrac{N-1}{N}$, so that the number of possible choices of $\mathcal{G}_{train}\subseteq A_{ij}$ of cardinality $pN^{n}$ is given by $\tfrac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$. Now we bound the probability of the event on the left hand side in (4) via a union bound over the $Nn$ possible choices of $A_{ij}$ for $1\leq i\leq n,1\leq j\leq N$, so that
$\mathbb{P}\left\{\begin{array}[]{l}\exists i\leq n,\ \exists j\leq N,\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train},\\ G_{i}\neq j\end{array}\right\}$ $\leq$ $nN\frac{(pN^{n})!((1-p)N^{n})!}{(N^{n})!}\frac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$
$=$ $nN\frac{N^{n}\frac{N-1}{N}}{N^{n}}\frac{N^{n}\frac{N-1}{N}-1}{N^{n}-1}\cdots\frac{N^{n}\frac{N-1}{N}-pN^{n}+1}{N^{n}-pN^{n}+1}$
$\leq$ $nN\left(1-\frac{1}{N}\right)^{pN^{n}},$
in which in the first line we used a union bound together with the computations of combinatorial coefficients from above, in the second line we performed a simplification and reordering, and in the third line we used the fact that $1-1/N=(N^{n}(N-1)/N)/N^{n}>(N^{n}(N-1)/N-i)/(N^{n}-i)$ for all $1\leq i<pN^{n}-1$.
Finally, considering the limit $N^{n}\to\infty$ of the upper bound in (4), we have two cases: (1) if $N\geq 2$ stays bounded and $n\to\infty$, then we can bound $nN(1-1/N)^{pN^{n}}\leq C_{1}nC_{2}^{C_{3}^{n}}$ with constants $C_{2}\in(0,1)$ and $C_{3}>1$, and by taking logarithms we check that the quantity tends to zero; (2) if $N\to\infty$ and $n\geq 2$, and in this case $(1-1/N)^{pN^{n}}\sim e^{-pN^{n-1}}$ and again by taking logarithms we find that the quantity tends to zero. ∎
In fact Prop. D.2 is gives us guarantees that it will be possible to observe all options for each $G_{i}$ in the training set, however this may not be sufficient for the main aim of recovering the compositional structure of $\mathcal{G}$. However we next prove, in an extension of Prop. D.2, that actually we will also observe with high probability any possible combination of finitely many factors:
Theorem D.3.
Under the same hypotheses on $\mathcal{G},\mathsf{GenX},\mathcal{G}_{train},p$ as in Prop. D.2, for any $k\leq n$ we have the following.
$\mathbb{P}\left\{\left.\begin{array}[]{l}\exists i_{1}<\dots<i_{k}\leq n\\ \exists j_{1},\dots,j_{k}\leq N\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train}\end{array}\right.(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\right\}$ (9)
$\leq\frac{n!}{k!(n-k)!}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}},$ (10)
which for fixed $p,k$ tends to zero as $N^{n}\to\infty$.
Proof.
As mentioned before, roughly the overall proof strategy is the same as for Prop. D.2. We introduce the sets $A^{j_{1},\dots,j_{k}}_{i_{1},\dots,i_{k}}=A^{\mathbf{j}}_{\mathbf{i}}$ and observe that the number of choices of $i_{1}<\cdots<i_{k}\leq n$ is $\tfrac{n!}{k!(n-k)!}$ while the number of choices of $j_{1},\dots,j_{k}\leq N$ is $N^{k}$. Thus the number of sets of the form $A^{\mathbf{j}}_{\mathbf{i}}$ is $N_{A}:=\tfrac{N^{k}n!}{k!(n-k)!}$. Furthermore, each set $A^{\mathbf{j}}_{\mathbf{i}}$ only imposes the constraint of entries of indices $i_{1},\dots,i_{k}$ from a $\mathbb{G}$ to avoid one combination of values $\vec{j}$, and thus it has cardinality $|A^{\mathbf{j}}_{\mathbf{i}}|=N^{n-k}(N^{k}-1)$. Then the number of choices of $\mathcal{G}_{train}\subseteq A^{\mathbf{j}}_{\mathbf{i}}$ of cardinality $pN^{n}$ is just the combinatorial coefficient $\tfrac{(N^{n-k}(N^{k}-1))!}{(pN^{n})!(N^{n-k}(N^{k}-1-pN^{k}))!}$ and by union bound we then find by the same reasoning as in Prop. D.2,
[…]
$=$ $N_{A}\frac{N^{n}-N^{n-k}}{N^{n}}\frac{N^{n}-N^{n-k}-1}{N^{n}-1}\cdots\frac{N^{n}-N^{n-k}-pN^{n}+1}{N^{n}-pN^{n}+1}$
$\leq$ $N_{A}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}=\frac{n!}{k!(n-k)!}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}.$
This proves (9). In order to show that the bound tends to zero as $N^{n}\to\infty$ at fixed $p>0,k\leq n$, we first note that $\tfrac{n!}{k!(n-k)!}\leq n^{k}/k!$ and thus (as $k$ is assumed fixed) it suffices to show that
$n^{k}N^{k}\left(1-\frac{1}{N^{k}}\right)^{pN^{n}}\to 0\quad\text{as}\quad N^{n}\to\infty.$
Here again, we can proceed with the discussion of the two cases ($N$ bounded and $n\to\infty$, or $N\to\infty$) exactly as in the end of the proof of Prop. D.2, to conclude. ∎
Remark D.4 (Extension of Prop. D.2 and Theorems D.3 and D.1 to general $\mathcal{G}$).
If, for a choice of $n\geq 2$ and $N_{1},\dots,N_{n}\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N_{i}\}$, then $|\mathcal{G}|=N_{1}\cdots N_{n}$ and by carefully following the same steps as in the proof of Prop. D.2 we find
$\mathbb{P}\left\{\begin{array}[]{c}\exists i\leq n,\quad\exists j\leq N_{i},\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train},\\ \text{it holds }G_{i}\neq j\end{array}\right\}\leq\sum_{i=1}^{n}N_{i}\left(1-\frac{1}{N}_{i}\right)^{pN_{1}\dots N_{n}},$ (11)
Again, this tends to zero for fixed $p$ as $|\mathcal{G}|=N_{1}\dots N_{n}\to\infty$: to prove it, we can proceed as follows. In a first case, if $n>2$ stays bounded, then we can discuss separately for each $1\leq i\leq n$ whether $N_{i}\to\infty$ or whether it stays bounded, as done in the proof of Prop. D.2, and in either case the corresponding summand on the left of (4) tends to zero, as in case (2) from the proof of Prop. D.2. Otherwise, $n\to\infty$ and we can proceed analogously to case (1) from the end of proof of Prop. D.2.
For the extension of Thm. D.3 to general $\mathcal{G}$ we get the bound
$\mathbb{P}\left\{\left.\begin{array}[]{l}\exists i_{1}<\dots<i_{k}\leq n\\ \exists j_{1},\dots,j_{k}\leq N\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train}\end{array}\right.(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\right\}$ (15)
$\leq\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}N_{i_{1}}\cdots N_{i_{k}}\left(1-\frac{1}{N_{i_{1}}\cdots N_{i_{k}}}\right)^{pN_{1}\cdots N_{n}},$ (16)
and the strategy for its proof and for the proof that this goes to zero if $k,p$ fixed and $|\mathcal{G}|=N_{1}\cdots N_{n}\to\infty$ are a direct extension of the above.
Theorem D.1 is a direct consequence of assumption (A) and Theorem D.3, and thus it also directly extends to the general case.
Appendix E Proof that final state reconstruction allows for wider message choice freedom
We first recall the setting. Consider two fixed maps
$R^{\star}:V^{C}\to\mathcal{X},\quad S^{\star}:\mathcal{D}\subseteq\mathcal{X}\to V^{C},$
and let $\mathcal{R}_{[i]}:V^{C}\to V^{i}$ be the restriction to the first $i$ tokens, i.e. in previous notation, for $m=(m_{1},\dots,m_{C})\in V^{C}$ we set $\mathcal{R}_{[i]}(m):=m_{[i]}=(m_{1},\dots,m_{i})$.
For $R^{\star},S^{\star}$ as above and $S:\mathcal{D}\subseteq\mathcal{X}\to V^{C}$ we set, for $x\in\mathcal{D}$
$d_{i}^{S}(x):=d_{\mathcal{X}}(R^{\star}(S^{\star}(x)_{[i]})),R^{\star}(S(x)_{[i]}))=d_{\mathcal{X}}(R^{\star}(\mathcal{R}_{[i]}(S^{\star}(x)))),R^{\star}(\mathcal{R}_{[i]}(S(x)))).$
We also defined the following spaces:
$\mathcal{S}_{full}^{\epsilon}:=\Big\{S:\ \mathbb{E}_{x}\Big[\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x)\Big]<\epsilon\Big\},\quad\mathcal{S}_{final}^{\epsilon}:=\left\{S:\ \mathbb{E}_{x}\left[d_{C}^{S}(x)\right]<\epsilon\right\}.$ (17)
Our main result here is the following:
Proposition E.1.
With the above notations, we have the following
1. For $i<C$ and any choice of $M,\epsilon>0$, the restrictions on $i$-token sub-messages of $S(x)$ by requiring $S\in\mathcal{S}_{full}^{M}$ are more restrictive than those imposed by requiring $S\in\mathcal{S}_{final}^{\epsilon}$, i.e. we have $\mathcal{R}_{[i]}(\mathcal{S}_{full}^{M})\subset\mathcal{R}_{[i]}(\mathcal{S}_{final}^{\epsilon})$.
2. For the full message case $i=C$ in general we have $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{C\epsilon}$. If we further assume that $\mathbb{E}_{x}[d_{j}^{S}(x)]$ is non-increasing in $j$ then we have the stronger inclusion $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon}$.
Proof.
Item 1 is direct, as no restriction on intermediate sub-messages $S(x)_{[i]}=\mathcal{R}_{[i]}(S(x))$ by the requirement $S\in\mathcal{S}_{final}^{\epsilon}$.
For item 2, we observe that
$\frac{1}{C}d_{C}^{S}(x)\leq\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x),$
and by taking expectation we find that $\mathbb{E}_{x}\left[\frac{1}{C}\sum_{i}d_{i}^{S}(x)\right]<\epsilon$ implies $\mathbb{E}_{x}[d_{C}^{S}(x)]\leq C\epsilon$, and thus $\mathcal{S}^{\epsilon}_{full}\subseteq\mathcal{S}^{C\epsilon}_{final}$, which gives the first part of the statement.
For the second part of item 2 we note that if $\mathbb{E}_{x}[d_{i}^{S}(x)]\geq\mathbb{E}_{x}[d_{C}^{S}(x)]$ for all $i\leq C$ we get
$\mathbb{E}_{x}\left[d_{C}^{S}(x)\right]\leq\min_{1\leq j\leq C}\mathbb{E}_{x}[d_{j}^{C}(x)]\leq\frac{1}{C}\sum_{i=1}^{C}\mathbb{E}_{x}\left[d_{i}^{S}(x)\right]=\mathbb{E}_{x}\left[\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x)\right],$
and thus $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon}$ as desired. ∎
The freedom of exploration ensured by item 1 of the proposition is our main motivation: by only constraining the reconstruction of the full message, we allow much more freedom for the encoding strategy of $S(x)$ at earlier sub-messages at each iteration.
Item 2 of Prop. E.1 shows that the two losses in the definition of $\mathcal{S}_{full},\mathcal{S}_{final}$ give comparable guarantees. At the beginning of iterations we will need to use the general bound with $\epsilon/C$, and as $R^{\star}$ will have been trained on more and more messages during the iteration, the hypothesis of $\mathbb{E}_{x}[d_{i}^{S}(x)]$ non-increasing will become true allowing for the sharper bound with the same value of $\epsilon$, as $R^{\star}$ (an LSTM taking as input successive message tokens) will become efficient in using sub-messages of pairs $(m,m^{\prime})=(S(x),S^{\star}(x))$ to gradually distill information allowing to distinguish them for a large set of $S,x$.
Appendix F Regularization to counteract Holistic Encodings
In this section, we include rigorous results useful to justify the regularization of our Imitation phase, aimed at balancing compressibility and efficiency.
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. [Kirby2002, KIRBY201587]). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
Our aim in this section is threefold:
  1. 1.
    In §F.1 we set rigorous definitions of Holistic Encodings within the setting of §2.1 and §$2.$
  2. 2.
    In §F.2 we check that compressibility is required for lowest Kolmogorov Complexity encodings.
  3. 3.
    We verify that optimizing for just expressivity and efficiency on our training set, will produce Holistic Encodings, and thus is not satisfactory for reconstructing $\mathcal{G}$ (see §F.3);
  4. 4.
    We give a few possible options on how to enrich the loss functions in our IL setup, aimed at establishing a better balance between efficiency and compressibility (see §F.4).
F.1 Definitions of Holistic and Compositional encodings, and role of the compressibility condition
In the classical IL setup [Kirby2002, KIRBY201587], a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but disregards the compositional structure of the "language" to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
Definition F.1.
In the setting from §2.1 and §2.2, sender $S:\mathcal{X}\to\mathcal{M}$ determines a holistic encoding if it satisfies expressivity but the messages associated to data generated as $\mathsf{GenX}(\mathbf{G})$ do not have a compositional structure in terms of generating factors $\mathbf{G}$, as specified by Definition F.2 below.
Note that, as in [Kirby2002, KIRBY201587], Definition F.1 of holistic encodings refers to "respecting compositional structure", which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
Definition F.2.
In the setting of §2.1 and §2.2, assume that generating factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ have $G_{i}\in[1:N_{i}]$ for some $N_{1},\dots,N_{n}\geq 2$. Then sender $S:\mathcal{X}\to\mathcal{M}=V^{C}$ will be defined to respect the structure of generating factors $\mathbf{G}$ if for each $1\leq i\leq n$ there exist an injective mapping $E_{i}:[1:N_{i}]\to V^{C_{i}}$, for which
$\forall\mathbf{G}\in\mathcal{G},\ S(\mathsf{GenX}(\mathbf{G}))=E_{1}(G_{1})\cdots E_{n}(G_{n})\quad\text{ and }\quad\sum_{i=1}^{n}C_{i}=C.$
The compressibility condition on the sender-receiver protocol just requires that the protocol be not holistic, and thus it is fully specified via definitions F.1 and F.2. Furthermore, Definition F.2 fully specifies in which way the map $\mathcal{G}\to\mathcal{M}$ successfully approximates an isomorphic reconstruction of $\mathcal{G}$, if it satisfies conditions compressibility and expressivity.
F.2 Encodings optimizing for only expressivity and efficiency have Kolmogorov Complexity much higher than compositional ones
By the classical learning theory principle, if our sender generates encodings with lower complexity, these should be easier to learn by the receiver, in the sense that the receiver will tend to generalize with more ease form good accuracy on the training set to similar accuracy on the test set.
As a manageable measure of complexity we will use Kolmogorov Complexity (KC), and we verify that in our setting, encodings satisfying expressivity and efficiency will necessarily have higher KC than compositional ones. In particular, the receiver will have more trouble reconstructing from such encodings, justifying our push to enforce the compressibility condition.
Encodings of $\mathcal{D}_{train}$ satisfying expressivity and efficiency include all injective maps
$\mathsf{enc}:\mathcal{D}_{train}\to V^{C_{\star}},\quad C_{\star}=\lceil\log_{|V|}|\mathcal{D}_{train}|\rceil,$ (18)
in which $\lceil a\rceil$ denotes the smallest integer larger than real number $a$. For a random map $\mathsf{enc}$ as above, the encoding to be Kolmogorov-irreducible, so that the expected KC of a random $\mathsf{enc}$ would be up to constant factor the one corresponding to explicitly enumerating the $|\mathcal{D}_{train}|$ strings of length $C_{\star}$ in alphabet $V$ that describe the encodings of each element of the training set. Each such string thus requires $\approx C_{\star}\log_{2}|V|$ bits. Thus if $U$ is the uniform distribution over injective maps (18) then we have
$\mathbb{E}_{\mathsf{enc}\sim U}[KC(\mathsf{enc})]\approx|\mathcal{D}_{train}|C_{\star}\log_{2}|V|\approx|\mathcal{D}_{train}|\log_{2}|\mathcal{D}_{train}|.$ (19)
On the other hand, to specify a compositional encoding $\mathsf{comp}$ as in Definition F.2 requires only assigning a string of length $C_{i}$ in alphabet $V$ for every generating factor $G_{i}$, where we can take $C_{i}=\lceil\log_{|V|}N_{i}\rceil$. Thus such string requires $\approx\log_{2}N_{i}$ bits, and the whole encoding has KC given by
$KC(\mathsf{comp})\approx\sum_{i=1}^{n}\log_{2}N_{i}=\log_{2}|\mathcal{D}_{train}|.$ (20)
If we assume that $|\mathcal{D}_{train}|/|\mathcal{D}|=p\in(0,1)$ and we take $|\mathcal{D}|\to\infty$, then we get
[…]
The computation (21) shows that
Proposition F.3.
Under the same hypotheses on $\mathcal{G}$ as in the previous section, consider the limit $|\mathcal{D}|\to\infty$ and assume that $p=|\mathcal{D}_{train}|/|\mathcal{D}|$ stays bounded away from zero in the limit. Then the expected Kolmogorov complexity of encodings satisfying expressivity and efficiency becomes overwhelmingly larger than the one of compositional encodings:
$\frac{\mathbb{E}_{\mathsf{enc}\sim U}[KC(\mathsf{enc})]}{KC(\mathsf{comp})}\sim p|\mathcal{D}|\to\infty.$
F.3 Encodings that optimize efficiency under expressivity over $\mathcal{D}_{train}$ are holistic
First, we clarify the definitions:
  • The requirement of "expressivity over $\mathcal{D}_{train}$" means that the encodings are injective over $\mathcal{D}_{train}$.
  • The requirement of "efficiency over $\mathcal{D}_{train}$" means that the encoding maps $\mathcal{D}_{train}$ into $V^{C_{\star}}$ for the minimum value of $C_{\star}$.
Similarly to previous subsection (see (18)), keeping in mind that $|\mathcal{D}_{train}|=p|\mathcal{D}|=p|\mathcal{G}|$, with fixed $p\in(0,1)$, we have
$C_{\star}=\lceil\log_{|V|}|\mathcal{D}_{train}|\rceil=\lceil\log_{|V|}p|\mathcal{G}|\rceil\leq\lceil\log_{|V|}|\mathcal{G}|\rceil-\lfloor|\log_{|V|}p|\rfloor.$ (22)
For compositional encodings of $\mathbf{G}=[G_{1},\dots,G_{n}]$ the encoding will require to set apart separate message sub-strings of length $C_{i}=\lceil\log_{|V|}N_{i}\rceil$ for factor $G_{i}$ (see Def. F.2), thus the required length for the encoding will be
$C_{0}:=\sum_{i=1}^{n}\lceil\log_{|V|}N_{i}\rceil\stackrel{{\scriptstyle(\star)}}{{\geq}}\left\lceil\sum_{i=1}^{n}\log_{|V|}N_{i}\right\rceil=\lceil\log_{|V|}|\mathcal{G}|\rceil\stackrel{{\scriptstyle(\star\star)}}{{\geq}}C_{\star}+\lfloor|\log_{|V|}p|\rfloor,$ (23)
in which $(\star\star)$ follows from (22).
We now discuss when the inequality signs in (23) are sharp or not, since sharp inequality $C_{0}>C_{\star}$ implies our claim that optimizing for efficiency under expressivity implies non-compositional (i.e., holistic) encodings:
  1. 1.
    If we use relatively small $|V|\lesssim 10$ (assuming commonly used values $p\lesssim 0.1$), we will have
    $|V|\leq\frac{1}{p},$ (24)
    and thus $\log_{|V|}p\leq-1$, and thus the second inequality $(\star\star)$ in (23) is guaranteed to be sharp, showing the desired strict inequality $C_{0}>C_{\star}$ independently of the sizes of factor ranges $N_{i},i=1,\dots,n$.
  2. 2.
    If the gaps between $\log_{|V|}N_{i}$ and the lowest integer larger or equal to it, sum to a value of at least $1$, i.e.,
    $\sum_{i=1}^{n}\left(\lceil\log_{|V|}N_{i}\rceil-\log_{|V|}N_{i}\right)\geq 1,$ (25)
    then $(\star)$ becomes sharp, again guaranteeing $C_{0}>C_{\star}$ as desired. This could be probable to happen in our setting, and becomes more likely for large $|V|>1/p$ for which (24) fails, given that the $N_{i}$ are supposed to be unknown, so it is likely that for large $|V|>1/p$ for a few of the $N_{i}$ the gaps in (25) are nontrivial. For example, (25) holds true for $|V|>1/p\geq 10$ if, say, at least two of the $G_{i}$ are binary factors, so that $N_{i}=2$.
In summary we have the following:
Proposition F.4.
Under the same running assumptions over $\mathcal{G}$, $p$, we have that optimizing only for efficiency of encodings efficiency under injectivity constraints expressivity over $\mathcal{D}_{train}$ is guaranteed to produce holistic (i.e., non-compositional) encodings if one or both of the conditions (24) and (25) holds.
F.4 Possible choices of encoded message regularization
As seen above, if we just optimize for efficiency and expressivity then we are likely to get holistic encodings (Prop. F.4), which will then make it hard for the receiver to generalize the decoding strategy outside the training set due to having overwhelmingly higher expected complexity (Prop. F.3).
This proves that it is important to incentivize condition compressibility, i.e., to push encoding strategies away from overly efficient holistic encodings. In this section we discuss a few approaches to do this in practice, explaining our choice of compressibility regularization.
  1. 1.
    Entropy maximization.
    Background and standard approach. Introduced by [entropymax] for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
    $h(\xi,W,x)=-\ln Pr(y=\xi|W,x),$ (26)
    such that if we take the expectation over $\xi$ we get:
    $\mathbb{E}[h(\xi,W,x)|W,x]=-\sum_{\xi}Pr(y=\xi|W,x)\ln Pr(y=\xi|W,x),$ (27)
    which is the entropy of the network. Therefore $h$ is an unbiased estimator of the entropy. They also note that, if $w_{i}$ and $x_{i}$ with $i\in U$ represent the weights and pattern of single neurons in a feed forward network, $\xi_{i}$ the $i_{th}$ position in a $n_{u}$-tuple $\xi$, and let $g_{i}$ be the probability density describing $y_{i}$, i.e:
    $g_{i}(\xi,w^{i},x^{i})=Pr(y_{i}=\xi|w^{i},x^{i}),$ (28)
    then, for any feed forward network,
    $Pr(y=\xi|W,x)=\prod_{i\in U}g_{i}(\xi_{i},w^{i},x^{i}),$ (29)
    and therefore
    $\ln Pr(y=\xi|W,x)=-\sum_{i\in U}g_{i}(\xi_{i},w^{i},x^{i})=h(y,W,X).$ (30)
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in [ren2020Compositional].
    Naive adaptation to our setting is not practical. We take $x\in\mathcal{X}$ and $\xi\in\mathcal{M}=V^{C}$ and recall that in our notation $m_{t}$ indicates the $t$-th token of message $m$, and $m_{[t]}$ is the notation for the initial segment $m_{1},m_{2},\dots,m_{t}$, so that in particular $m_{[C]}=m$.
    Importantly, in our setting, to avoid high variance in estimation due to the large message space (see §G), we approximate the probability distribution of the RNN-produced message token at position $t\leq C$, denoted here $p(m_{t}|x,m_{[t-1]})$, as a $\delta$ distribution over the vocabulary $V$ by using the Gumbel-softmax trick:
[…]
    and therefore:
    $Pr(m=\xi|W,x)=\prod_{t=1}^{C}Pr(m_{t}=\xi_{t}|W,x,m_{[t-1]}).$ (32)
    Note that due to (31), for all but a single one of the $|V|^{C}$ possible values of $\xi$ the quantity (32) is $\approx 0$, and thus $h$ becomes a poor estimator for the entropy of $p(m_{t}|x,m_{[t-1]})$, as we get
    $h(m,W,x)=-\ln\sum_{t=1}^{C}Pr(m_{t}=\xi_{t}|x,W,m_{[t-1]})\approx\infty.$ (33)
    In particular, it becomes clear that optimizing this $h$ becomes unfeasible with gradient methods.
    Our proposal for regularization. Our model can be seen as a different approximation of the entropy of the message. Instead of using the policy weights as probability values, we use the law of large numbers to estimate the probability of each message. For a batch $\mathcal{B}\subset\mathcal{X}$ consisting of $N_{batch}$ training examples and for given sender encoding protocol $S_{\phi}$:
    $\text{For a sufficiently large $N_{batch}$:}\quad\Pr(m=\xi\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)=\xi\}|}{N_{batch}}\\$ (34)
    Yet again, due to the exponential size of the message space, utilizing this probability to estimate the entropy is still not practical. We will assume approximate independence between symbol positions, which allows to get the stronger approximation:
    $\Pr(m_{t}=v\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)_{t}=v\}|}{N_{batch}},$ (35)
    as well as the additivity of entropy to be applied to (27):
    $H(Pr(m|x,\phi))$ $:=$ $-\sum_{\xi\in V^{C}}Pr(m=\xi\mid x,\phi)\ln Pr(m=\xi\mid x,\phi)$ (36)
    $=$ $-\sum_{t=1}^{C}\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\ln Pr(m_{t}=v\mid x,\phi).$
    Then we can perform the following computations, in which we let $m,m^{\prime}$ be two i.i.d. copies of $m$:
[…]
    $\overset{\begin{subarray}{c}\text{($-x\ln x\geq x-x^{2}$}\\ \text{for $x\leq 1$)}\end{subarray}}{\geq}\sum_{t=1}^{C}\sum_{v\in V}\left[Pr(m_{t}=v\mid x,\phi)-Pr(m_{t}=v\mid x,\phi)^{2}\right]$
    $\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)\right$.
    $\quad\left.-\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)^{2}\right)$
    $=\sum_{t=1}^{C}\sum_{v\neq v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)$
    $=\sum_{t=1}^{C}Pr(m_{t}\neq m_{t}^{\prime})=\mathbb{E}[d_{H}(m,m^{\prime})]$
[…]
    In the above step $(\star)$ we used that
    $\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)$ $=$ $1=\left(\sum_{v\in V}Pr(m_{t}=v\mid x,\phi)\right)^{2}$
    $=$ $\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi).$
    The consequence of bound (1) is that maximizing the pairwise Hamming distance between predicted messages in a batch also maximizes a lower bound to the approximate entropy of the message space, justifying the choice of optimizing this quantity (which is easier to i include in practice) for favoring compressibility.
  2. 2.
    Contrastive Learning losses. In our framework, PDM can be seen as a form of contrastive loss, which we now show how to connect to standard contranstive losses from previous works.
    Contrastive loss functions such as the triplet loss [triplet] and NT-Xtent [chen2020simpleframeworkcontrastivelearning] work by defining positive and negative image pairs (usually through image augmentation), and aiming to minimize the distance between embeddings of positive pairs and maximize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair of samples for each $x$ in a mini-batch $N$ by augmenting each image. Each image in the batch therefore has 2N12N-1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
    $\ell_{i,j}=-\log\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum^{2N}_{k}\exp(sim(z_{i},z_{k})/\tau)}$ (38)
    For simplicity, let $\tau=1$ and thus
    $\ell_{i,j}=-sim(z_{i},z_{j})+\log\sum^{2N}_{k}\exp(sim(z_{i},z_{k})).$ (39)
    Here, we apply the same ideas but take this loss over an unaugmented batch, i.e the only application of this loss is for $z_{i}$ against itself. In this case it reduces to:
    $\ell_{i}=-1+\log\sum^{N}_{k}\exp(sim(z_{i},z_{k}))$ (40)
    In our representations, the embedding vectors from a batch $z_{i},i=1,\dots,N$ are $V\times C$ matrices whose columns are approximately one-hot vectors that represent the tokens of a message. For one-hot matrices $z_{i},z_{k}$ of this form, the dot product $z_{i}^{T}z_{k}$ is equivalent to counting the amount of matching rows of $z_{i},z_{k}$, i.e, the un-normalized Hamming distance of the associated message strings $m_{i},m_{k}\in V^{C}$. Since strings have length $C$ and $\|z_{i}\|^{2}=z_{i}^{T}z_{i}\simeq C$, we thus get:
    $sim(z_{i},z_{k})$ $=$ $tr(z_{i}^{T}z_{k})/\|z_{i}\|\|z_{k}\|\simeq\frac{1}{C}\sum_{l=1}^{C}\mathds{1}[\arg\max_{j}((z_{i})_{l,j})=\arg\max_{j}((z_{k})_{l,j})]$
    $\simeq$ $1-d_{H}(m_{i},m_{k}),$
    in which $d_{H}$ is the (normalized by $1/C$) Hamming distance. We can replace this expression in the unaugmented NT-Xtent loss:
    $\ell_{i}$ $=$ $-1+\log\sum_{k=1}^{N}\exp(1-d_{H}(m_{i},m_{k}))$
    $=$ $-1+N-\log\sum_{k=1}^{N}\exp(d_{H}(m_{i},m_{k}))$
    $\overset{\begin{subarray}{c}\text{(Jensen}\\ \text{ inequality)}\end{subarray}}{\leq}$ $-1+N-\log N-\frac{1}{N}\sum_{k=1}^{N}d_{H}(m_{i},m_{k}).$
    Then, taking the mean loss over the full batch we get:
    $\frac{1}{N}\sum^{N}_{i=1}\ell_{i}\leq-1+N-\log N-\frac{1}{N^{2}}\sum^{N}_{i=1}\sum^{N}_{k=1}d_{H}(m_{i},m_{k}).$ (41)
    We then include in the loss only the last term, i.e. the average of $d_{H}(m_{i},m_{k})$, which is the only term that actually depends on the batch elements, which forms an approximate upper bound on the NT-Xtent loss.
    Note that our approximation of $sim(z_{i},z_{k})$ by $1-d_{H}(m_{i},m_{k})$, on which the above is based, becomes sharp when the $z_{i}$ have actual one-hot entries, thus the approximation becomes better as our model has high accuracy. Furthermore, Jensen’s inequality, responsbile for the inequality between the term $\ell_{i}$ from NT-Xtent and our metric, becomes closer to equality messages $m_{1},\dots,m_{N}$ become uniformly spread across $V^{C}$, which is the case for minimizers of our metric as well. This means that, at least heuristically speaking, we can consider our metric as having similar minimizers to the NT-Xtent regularization.
    A precise matching lower bound to (41) is left for future work.
Appendix G Justification of our imitation step starting from classical probabilistic reconstruction games
We start by recalling classical probabilistic reconstruction games, after which we adapt the formulation to the setting of generative tasks, and we show that with a few justified simplifications we arrive to our formulation from the main paper.
G.1 Classical probabilistic reconstruction games
Recall that in reconstruction games, the two agent types are a teacher and student (or speaker and listener, or sender and receiver), parameterized by respectively $\theta,\phi$. The game consists in the teacher observing a signal $x$ distributed according to a probability distribution $p$, from which it produces a randomized message $m$ whose distribution conditional on $x$ is obtained according to policy $\pi_{\theta}(\cdot|x)$. The student or listener, observes message $m$ and produces a reconstructed signal distributed according to a conditional probability $\rho_{\phi}(\cdot|m)$. The game dynamics can be modeled by stipulating that together, teacher and student minimize the log-likelihood loss function
$\mathcal{L}_{\theta,\phi}=-\mathbb{E}_{x\sim p,m\sim\pi_{\theta}(\cdot|x)}\left[\log\rho_{\phi}(x|m)\right]$ (42)
in which $x$ represents the signal and $m$ the message used to reconstruct it, the speaker produces a reinforcement learning policy $\pi_{\theta}(\cdot|x)$ and the listener produces a reconstruction function $\rho_{\phi}(x|m)$.
G.2 Reconstruction games in generative tasks
While the above formulation (42) is formalized using the log-likelihood between distributions, for our generative task, the reconstruction modeled slightly differently: the task is to optimize, with high probability, the reconstruction error, measured by a distance between the signal $x$ and its reconstruction $x^{\prime}$, denoted $d(x,x^{\prime})$. Here the notation for $x,x^{\prime}$ and $d(x,x^{\prime})$ can be (ab)used to represent two alternative settings: either (a) $x,x^{\prime}$ are vectors in the autoencoder latent space and $d(x,x^{\prime})$ is a discrepancy (e.g. $d(x,x^{\prime})=|x-x^{\prime}|^{p}$) between them or (b) $x,x^{\prime}$ are reconstructed images and $d(x,x^{\prime})$ is a discrepancy (e.g. based on some ad-hoc norm or otherwise) between them. Thus we have the probabilistic error function
$\mathcal{L}_{int}^{\mathbb{P}}(\omega,\varphi):=\mathbb{E}_{x\sim p,m\sim\pi_{\varphi}(\cdot|x),x^{\prime}\sim\rho_{\omega}(\cdot|m)}d(x,x^{\prime}).$ (43)
The loss (43) is akin to a $d$-based Wasserstein distance analogue of (42).
G.3 The problem reduces to the case of a deterministic student
Note that when $d(x,x)=|x-x^{\prime}|^{2}$ for some Euclidean norm (or more generally, when $d$ is a convex function of such norm), then the minimization of (43) over distributions $\rho_{\omega}$ at fixed $m,\varphi$ must be achieved by
$x_{m}^{\varphi}\in\mathrm{argmin}_{x^{\prime}}\mathbb{E}_{x\sim p_{\varphi}(\cdot|m)}d(x,x^{\prime}),$ (44)
where if the set of $x$ is finite then $p_{\varphi}$ is the probability defined by
$p_{\varphi}(x^{\prime}|m)=\frac{p(x^{\prime})\pi_{\varphi}(m|x^{\prime})}{\sum_{y}p(y)\pi_{\varphi}(m|y)},$ (45)
and an analogous expression with integrals replacing the sum holds in general.
Under the above convexity hypotheses, the choice of $x_{m}^{\varphi}$ in (44) is unique, and for $d(x,x^{\prime})=|x-x^{\prime}|^{2}$ it coincides with the barycenter of $p_{\varphi}$. In this case, $\rho_{\omega}(\cdot|m)$ can be taken to be the Dirac distribution concentrated at $x_{m}^{\varphi}$, and optimization of (43) can be achieved by a deterministic reconstruction function $R_{\omega}(m)$, whose optimum value for given $\varphi$ will be $x_{m}^{\varphi}$ from (44). When reducing to optimizing only amongst deterministic $R_{\omega}$, the loss (43) rewrites in the simplified form
$\mathcal{L}_{int}(\omega,\varphi)=\mathbb{E}_{x\sim p,m\sim\pi_{\varphi}(\cdot|x)}d(x,R_{\omega}(m)).$ (46)
G.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in [havrylov2017emergence] and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE [williams1992simple], however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick [jang2017categoricalreparameterizationgumbelsoftmax], in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (46) as
$\mathcal{L}_{int}(\omega,\phi)=\mathbb{E}_{x\sim p}d(x,R_{\omega}(S_{\phi}(x))),$ (47)
which for $d(x,x^{\prime})=|x-x^{\prime}|$ is very similar to our actual interaction loss (1):
$\mathcal{L}_{int}(\omega,\phi)=\mathbb{E}_{x\in\mathcal{X}}\left[|x-R_{\omega}(S_{\phi}(x))|\right].$ (48)
Appendix H Example Generation: Reconstructions
The following reconstruction, Figure 5, examples were generated with the CELEBI model using PDM regularization with $\beta=1$ and $\lambda=1.5$. In each row, the leftmost image corresponds to the original input, while the subsequent images represent the predicted reconstructions $x_{i}=R_{\omega}(S_{\phi}(x)_{[1:i]})$ obtained from progressively longer message prefixes with the VAE backbone.
We observe that the reconstructions do not merely improve in pixel-level fidelity but rather exhibit semantic refinements across successive steps — variations in floor hue, object color, shape, or viewing angle. This progression suggests that the message space encodes the underlying generative factors $\mathcal{G}$ and supports a degree of compositional structure, where messages carry disentangled information.
Qualitative comparison across methods.
To further assess the representational properties of CELEBI, we qualitatively compare reconstructions across the different methods proposed in this work (see Figures 6 and 7). As illustrated in these figures, CELEBI reaches semantically faithful reconstructions significantly earlier in the decoding sequence than the baseline methods.
Moreover, the differences between reconstructions are more semantically pronounced than in baseline models, which often show subtler variations mainly related to texture or brightness. This qualitative evidence supports our hypothesis that CELEBI facilitates a more structured message space, enabling representations that are not only more compact but also more semantically coherent.
(a) Example 1
(b) Example 2
(c) Example 3
(d) Example 4
Figure 5: Progressive image reconstructions obtained at each decoding step by the receiver as it processes successive symbols of the message.
(a) Reconstruction using the baseline model.
(b) Reconstruction using partial decoding.
(c) Reconstruction using final state imitation.
(d) Reconstruction using PD+FiSI+PDM.
Figure 6: Reconstruction for the same example with different methods.
(a) Reconstruction using the baseline model.
(b) Reconstruction using partial decoding.
(c) Reconstruction using final state imitation.
(d) Reconstruction using PD+FiSI+PDM.
Figure 7: Reconstruction for the same example with different methods.
Appendix I Additional Language Metrics
Table 3: Comparison of proposed methods using language variation metrics from [conklin2023compositionality].
Synonymy $\downarrow$ Homonymy $\downarrow$ Freedom $\downarrow$ Entanglement $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.563\pm 0.026$ $0.695\pm 0.015$ $0.569\pm 0.026$ $0.831\pm 0.006$
Progressive Decoding (PD, ours) $0.640\pm 0.022$ $0.737\pm 0.013$ $0.646\pm 0.013$ $0.786\pm 0.007$
Imitation Message imitation (baseline) $0.601\pm 0.039$ $0.714\pm 0.027$ $0.613\pm 0.039$ $0.822\pm 0.010$
Full state imitation $0.597\pm 0.017$ $0.734\pm 0.021$ $0.604\pm 0.018$ $0.824\pm 0.007$
Final-State Imitation (FiSI, ours) $0.633\pm 0.038$ $0.751\pm 0.018$ $0.640\pm 0.038$ $0.842\pm 0.007$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.594\pm 0.027$ $0.739\pm 0.013$ $0.598\pm 0.027$ $0.805\pm 0.007$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.654\pm 0.013$ $0.753\pm 0.008$ $0.658\pm 0.013$ $0.780\pm 0.008$
(a) Shapes3D
Synonymy $\downarrow$ Homonymy $\downarrow$ Freedom $\downarrow$ Entanglement $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.595\pm 0.053$ $0.597\pm 0.061$ $0.597\pm 0.052$ $0.852\pm 0.021$
Progressive Decoding (PD, ours) $0.484\pm 0.044$ $0.486\pm 0.056$ $0.486\pm 0.044$ $0.786\pm 0.025$
Imitation Message imitation (baseline) $0.367\pm 0.034$ $0.393\pm 0.042$ $0.369\pm 0.034$ $0.837\pm 0.015$
Full state imitation $0.362\pm 0.045$ $0.424\pm 0.052$ $0.364\pm 0.044$ $0.832\pm 0.018$
Final-State Imitation (FiSI, ours) $0.324\pm 0.020$ $0.393\pm 0.018$ $0.328\pm 0.020$ $0.830\pm 0.008$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.365\pm 0.011$ $0.338\pm 0.019$ $0.367\pm 0.011$ $0.763\pm 0.005$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.347\pm 0.012$ $0.333\pm 0.010$ $0.349\pm 0.013$ $0.756\pm 0.005$
(b) MPI3D
In addition to the metrics presented in the main text of this work, we attempted to measure other metrics for qualitative linguistic variation. We were not able to find many such metrics in the literature, and thus restrict to the ones presented in [conklin2023compositionality]. Importantly, since this reference had a slightly different focus, variation in this context will not refer to the evolution of the language in time, but rather a departure from regularity, which masks an underlying compositional structure.
We study the four measures of variation presented in the paper, namely synonymy, homonymy, word order freedom and entanglement. Synonymy measures the presence of one-to-many mappings between atomic meanings and characters in a position, being minimized when each generating factor value is mapped to a single position and character. Homonymy measures the opposite, i.e., the presence of many-to-one mappings, and is minimized when each character in a position is mapped to a unique generating factor value. Word order freedom refers to the strictness of the mapping between generating factors and positions in the message (for example, if the mapping system always encodes the shape generating factor at the first position). It is minimized if all single generating factors are encoded in the same message position. Similarly, entanglement is minimized when all factors are encoded into unique positions in the message.
Important formular metricas, pero no encaja tan bien. Unico paper que conocemos. No clear patterns, but we argue it is because… An important assumption made by the authors for all these metrics is that meaning should be undivisably encoded into single positions of the message. We believe this scope of regularity is too narrow to capture compositionality-respecting mappings in out setting, as defined in F.2, and does not fit well when the space of messages is greatly larger than the amount of possible states. We found no clear trend or similar behavior across both tested datasets, save a small decrease in entanglement when using PD. This may suggest that the increased pressure for efficiency forces individual generating factors to be uniquely distributed in the message, however we believe there is not sufficient evidence to make such a claim considering the previously discussed pitfalls of the metrics.
To our knowledge there are no available official implementations for these metrics. Our versions can be found in the official repository for this work.
Appendix J Permutation Tests
J.1 Emergent language metrics
To assess the statistical significance of each component added to the Iterated Learning (IL) framework proposed in this work, we conducted permutation tests using the SciPy library [2020SciPy-NMeth]. The test statistic was defined as the difference in means, $\bar{A}-\bar{B}$, where $\bar{A}$ corresponds to the arithmetic mean of the evaluated mode and metric across 10 random seeds. The complete results are presented in Table 5(b).
Across both the Shapes3D and MPI3D datasets, we found that the inclusion of the PD module consistently reduced the useful message length, yielding a test statistic of approximately $2.5$ and a $p$-value below $0.05$. None of the other additions introduced in this work produced a statistically significant decrease in useful length relative to their respective baselines, supporting our hypothesis that PD imposes an efficiency pressure on the emergent language. Moreover, we observed a statistically significant increase in TopSim for the Shapes3D dataset, with a test statistic of $0.026$ and $p<0.05$.
Final-State Imitation had a statistically significant gain in TopSim over both baseline methods on MPI3D, with statistic of $\sim 0.02$ and $p<0.02$. PDM outperformed the KoLeo estimator in both reconstruction error and TopSim on Shapes3D, but achieved no statistically significant increase in MPI3D.
Table 4: Permutation tests for Shapes3D and MPI3D on emergent language metrics.
Metric Mode A Mode B Statistic p-Value
TopSim Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0263 0.0233
TopSim Full state imitation Final-State Imitation (FiSI; ours) -0.0267 0.1014
TopSim Full state imitation Message imitation (Baseline) -0.0018 0.9332
TopSim Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0249 0.2605
TopSim PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.0371 0.0004
Useful Length Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -2.5 0.0031
Useful Length Full state imitation Final-State Imitation (FiSI; ours) 0.0 1.0
Useful Length Full state imitation Message imitation (Baseline) 0.0 1.0
Useful Length Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0 1.0
Useful Length PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 1.3000 0.0996
Final MSE Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -0.0218 0.0657
Final MSE Full state imitation Final-State Imitation (FiSI; ours) -0.0130 0.4871
Final MSE Full state imitation Message imitation (Baseline) -0.0201 0.5679
Final MSE Final-State Imitation (FiSI; ours) Message imitation (Baseline) -0.0071 0.8384
Final MSE PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 0.0343 0.0431
(a) Shapes3D permutation test
Metric Mode A Mode B Statistic p-Value
TopSim Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0037 0.4416
TopSim Full state imitation Final-State Imitation (FiSI; ours) -0.0193 0.0019
TopSim Full state imitation Message imitation (Baseline) 0.0047 0.4395
TopSim Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0240 0.0007
TopSim PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.0060 0.4767
Useful Length Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -2.7000 0.0345
Useful Length Full state imitation Final-State Imitation (FiSI; ours) 0.0231 1.0
Useful Length Full state imitation Message imitation (Baseline) 0.2231 0.3998
Useful Length Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.2000 0.5820
Useful Length PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 0.2222 1.0
Final MSE Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0011 0.5894
Final MSE Full state imitation Final-State Imitation (FiSI; ours) -0.0018 0.1166
Final MSE Full state imitation Message imitation (Baseline) 0.0019 0.2870
Final MSE Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0037 0.0201
Final MSE PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.000012 0.9626
(b) MPI3D permutation test
J.2 Disentanglement permutation tests
In this section we evaluate the statistical significance of the differences between our model and the discrete baseline on the metrics used in table 2. Similar to the previous section, we conducted permutation tests, using the difference in means as a statistic for 10 random seeds. The complete results are presented in Table 5.
We found a significant difference in accuracy on the Shapes3D dataset, with a statistic of 0.114 and $p<0.01$. We found a significant descrease in RMSE on both datasets, with statistics of $0.227$ and $1.151$ for Shapes3D and MPI3D, respectively, and $p\leq 0.001$. We also observed a significant gain in disentanglement on both datasets, with statistics of $0.003$ and $0.023$ for Shapes3D and MPI3D, respectively, and $p<0.05$ for both.
Table 5: Permutation tests for Shapes3D and MPI3D on disentanglement, accuracy and RMSE using messages from discrete models as data for a linnear probe.
Metric Mode A Mode B Diff(A-B) p-Value
Accuracy VAE+EL VAE+CELEBI -0.1139 0.0018
RMSE VAE+EL VAE+CELEBI 0.2266 0.0010
DCI Disentanglement VAE+EL VAE+CELEBI -0.0034 0.0042
(a) Shapes3D permutation test
Metric Mode A Mode B Diff(A-B) p-Value
Accuracy VAE+EL VAE+CELEBI -0.0128 0.0570
RMSE VAE+EL VAE+CELEBI 1.1512 0.0010
DCI Disentanglement VAE+EL VAE+CELEBI -0.0230 0.0002
(b) MPI3D permutation test